### v1

Abstract
The main purpose of this work is to present a SIMD-vectorized implementation of the symplectic 16th-order 8-stage implicit Runge-Kutta integrator based on collocation with Gauss-Legendre nodes (IRKGL16-SIMD), and to show that it can outperform state-of-the-art symplectic explicit integrators for high-precision numerical integrations (in double-precision floating-point arithmetic) of non-stiff Hamiltonian ODE systems. Our IRKGL16-SIMD integrator leverages Single Instruction Multiple Data (SIMD) based parallelism (in a way that is transparent to the user) to significantly enhance the performance of the sequential IRKGL16 implementation. We present numerical experiments comparing IRKGL16-SIMD with state-of-the-art high-order explicit symplectic methods for the numerical integration of several Hamiltonian systems in double-precision floating-point arithmetic.
keywords:
Symplectic methods, Gauss implicit Runge-Kutta methods, SIMD-vectorization, non-stiff Hamiltonian ODE systems, Julia implementation
1 Introduction
The family of $s$-stage implicit Runge-Kutta schemes based on collocation with Gauss-Legendre nodes (IRKGL) are known to be symmetric, symplectic and super-convergent of order $2s$, and hence are very well suited for high precision long-term numerical integration of Hamiltonian systems [18, 11]. The computational cost per step of an implicit RK method is typically higher than that of an explicit RK method, as a system of algebraic equations must be solved at each step of implicit integrators. For non-stiff problems, fixed-point iterations are commonly employed to approximately solve such systems of algebraic equations at each step. A fixed-point implementation (based on the ideas presented in [1] and written in Julia language) of the $16$th order $8$-stage IRKGL scheme (IRKGL16) is included in the DifferentialEquations.jl suite [2, 16]. Our code, IRKGL16, includes a standard time-step adaptivity mechanism. However, for long-term integration of Hamiltonian systems, the use of a constant time step is recommended [18, 11, 13]. Therefore, in this work, we will focus on its application in constant step-size mode.
While there are no explicit symplectic Runge-Kutta schemes, optimized high-order explicit symplectic integrators are available for Hamiltonian systems with suitable structure. These integrators, that can be derived with the techniques of splitting and composition, are applicable when the Hamiltonian function can be decomposed into the sum of two, three, or more exactly solvable Hamiltonians [5, 11, 13].
It is widely acknowledged that high-order explicit symplectic methods are more efficient than IRKGL schemes for computations in double-precision floating-point arithmetic [11]. In fact, to achieve similar accuracy over a given integration interval, our sequential implementation of IRKGL16 usually requires more CPU time than the most effective high order explicit symplectic integrators in double-precision computations.
However, it is well established [11] that the efficiency of implicit Runge-Kutta methods can be significantly enhanced through various forms of parallelism. Single Instruction Multiple Data (SIMD), often referred to as SIMD-vectorization, is a type of parallelism that can be leveraged in most modern CPU cores, enabling vector computations at a cost comparable to scalar operations. The structure of the IRKGL16 integrator is particularly well-suited for exploiting SIMD parallelism with vectors containing eight (or alternatively, four or two) floating-point numbers.
The primary objective of this work is to present a SIMD-vectorized implementation of IRKGL16 in constant step-size mode (referred to as IRKGL16-SIMD) and to demonstrate its superiority over state-of-the-art explicit symplectic integrators for high-precision numerical integrations in double-precision floating-point arithmetic for several relevant Hamiltonian problems with a separable structure.
Our implementation makes use of the Julia package SIMD [19] to explicitly handle computations on vectors containing a fixed number of floating-point numbers. The algorithm IRKGL16-SIMD automatically performs vectorization, ensuring that the process remains transparent to the user.
The integrator IRKGL16-SIMD is designed to numerically integrate non-stiff systems of (non-necessarily Hamiltonian) ODEs whose right-hand side is implemented as a generic Julia function using arithmetic operations and elementary functions supported by the SIMD package.
The paper is organized as follows: Section 1 provides an introduction; Section 2 describes a vectorized fixed-point implementation of symplectic implicit Runge-Kutta (IRK) schemes; Section 3 presents implementation details of the IRKGL16-SIMD method; Section 4 reviews state-of-the-art symplectic explicit integrators; Section 5 reports numerical experiments; and Section 6 presents the conclusions.
2 Vectorized fixed-point implementation of symplectic IRK schemes
2.1 Implicit RK methods
We consider initial value problems for systems of ordinary differential equations (ODEs) of the form:
$\frac{d}{dt}y=f(t,y),\quad y(t_{0})=y_{0},$ (1)
where $f:\mathbb{R}^{D+1}\to\mathbb{R}^{D}$ is a sufficiently smooth function and $y_{0}\in\mathbb{R}^{D}$.
Runge-Kutta (RK) methods are widely used one-step integrators for numerically solving ODEs. These methods are specified by the coefficients in the Butcher tableau:
[…]
The RK approximations $y_{n}\approx y(t_{n})$ to the solution $y(t)$ at times $t_{n}=t_{0}+nh$ for $n=1,2,3,\dots$ are given by:
$y_{n}=y_{n-1}+h\sum_{i=1}^{s}b_{i}f(t_{n-1}+c_{i}h,Y_{n,i}),$ (2)
where
$Y_{n,i}=y_{n-1}+h\sum_{j=1}^{s}a_{ij}f(t_{n-1}+c_{j}h,Y_{n,j}),\quad i=1,\ldots,s.$ (3)
Here, each $Y_{n,i}$ is an approximation of the state vector $y(t)$ at the intermediate time $t=t_{n-1}+hc_{i}$ computed within each time-step. The coefficients $a_{ij}$, $b_{i}$, and $c_{i}$ from the Butcher tableau determine the specific RK method and its properties.
If the matrix $A=(a_{ij})$ is lower triangular the Runge-Kutta method is explicit. Otherwise, it is an implicit Runge-Kutta (IRK) scheme, and the stage vectors $Y_{n,i}$ at each step are defined implicitly by the equation (3).
2.2 IRK methods of collocation type
In a collocation method based on the nodes $c_{1},\ldots,c_{s}\in[0,1]$, the approximations $y_{n}\approx y(t_{n})$ to the solution $y(t)$ at times $t_{n}=t_{0}+nh$ are obtained as follows: for each $n=1,2,3,\dots$, consider the polynomial function $P_{n}:\mathbb{R}\to\mathbb{R}^{D}$ of degree $s$ uniquely determined by the following conditions,
  • $P_{n}(t_{n-1})=y_{n-1}$,
  • $P_{n}^{\prime}(t_{n-1}+c_{i}h)=f(t_{n-1}+c_{i}h,P_{n}(t_{n-1}+c_{i}h))$ for $i=1,\ldots,s$.
then, set $y_{n}=P_{n}(t_{n})$.
This is equivalent [11] to the application of the IRK method (2)–(3) with coefficients $b_{i},a_{ij}$ uniquely determined by the following conditions:
$\begin{split}\sum_{j=1}^{s}a_{ij}c_{j}^{k-1}&=\frac{c_{i}^{k}}{k},\quad i=1,\ldots,s,\quad k=1,2,\ldots,s,\\ \sum_{j=1}^{s}b_{j}c_{j}^{k-1}&=\frac{1}{k},\quad k=1,2,\ldots,s.\end{split}$ (4)
In the specific case where $c_{1},\ldots,c_{s}\in(0,1)$ are the shifted Gauss-Legendre nodes–defined by $c_{i}=(1+x_{i})/2$, where $x_{1},\ldots,x_{s}$ are the zeroes of the Legendre polynomial of degree ss–the resulting IRK scheme of collocation type, which we refer to as the $s$-stage IRKGL scheme, achieves an order of accuracy $2s$ [11].
2.3 Symplectic IRK methods with floating-point coefficients
We consider an autonomous Hamiltonian system of the form
$\frac{dy}{dt}=J^{-1}\nabla H(y),\quad y\in\mathbb{R}^{2d},$ (5)
where $H:\mathbb{R}^{2d}\to\mathbb{R}$ is the Hamiltonian function and $J$ is the standard symplectic matrix
$J=\begin{pmatrix}0&I_{d}\\ -I_{d}&0\end{pmatrix}$.
The exact flow of this system, denoted by $\varphi_{H}^{t}$, is symplectic, meaning that it preserves the symplectic form: for all $t$, the Jacobian matrix $D\varphi_{H}^{t}(y)$ satisfies
$\left(D\varphi_{H}^{t}(y)\right)^{T}JD\varphi_{H}^{t}(y)=J.$
It is therefore desirable that numerical integrators used to approximate such flows also preserve this structure, in order to ensure long-time stability and accurate qualitative behavior of the solution.
An implicit Runge-Kutta (IRK) method is symplectic if and only if [18] its coefficients satisfy the condition
$b_{i}a_{ij}+b_{j}a_{ji}=b_{i}b_{j},\quad 1\leqslant i,j\leqslant s.$ (6)
This condition ensures that the discrete flow defined by the IRK method is symplectic, and is satisfied in particular by the $s$-stage IRKGL schemes [18, 13].
As shown in [12], if the coefficients $b_{i},a_{ij}$ are replaced by machine numbers $\tilde{b}_{i},\tilde{a}_{ij}$, then the resulting implicit RK scheme typically fails to satisfy the simplectic condition (6). To ensure that the symplectic property is exactly preserved with double-precision floating-point coefficients, we reformulate the IRKGL schemes as follows [1]:
$Y_{n,i}$ $=y_{n-1}+\sum^{s}_{j=1}{\mu_{ij}\,L_{n,j},\quad L_{n,i}=hb_{i}f(t_{n-1}+hc_{i},Y_{n,i})},\quad i=1,\ldots,s,$ (7)
$y_{n}$ $=y_{n-1}+\sum^{s}_{i=1}L_{n,i},$ (8)
where $\mu_{ij}=a_{ij}/b_{j},\quad 1\leqslant i,j\leqslant s$. If the coefficients $b_{i}$, $c_{i}$, and $\mu_{ij}$ are replaced by the double-precision floating-point coefficients $\tilde{b}_{i}$, $\tilde{c}_{i}$, and $\tilde{\mu}_{ij}$, then the resulting scheme is still symplectic provided that $\tilde{\mu}_{ij}$ is set for $1\leqslant i<j\leqslant s$ as $1-\tilde{\mu}_{ji}$.
2.4 Vectorized fixed-point iteration
The primary challenge in implementing implicit Runge-Kutta (IRK) methods is the efficient solution of the nonlinear system for the internal stages. For non-stiff problems, fixed-point iteration is generally recommended. In collocation-type IRK methods, it is standard practice to initialize the fixed-point iteration with an estimate of the internal stages extrapolated from the previous time-step, that is,
$Y_{n,i}\approx P_{n-1}(t_{n-1}+hc_{i}),\quad i=1,\ldots,s.$ (9)
It is not difficult to check that
$P_{n-1}(t_{n-1}+hc_{i})=y_{n-1}+\sum_{j=1}^{s}\nu_{ij}L_{n-1,j},\quad i=1,\ldots,s,$ (10)
where the coefficients $\nu_{ij}$ are uniquely determined from the following equations
$\sum_{j=1}^{s}\nu_{ij}(c_{i}-1)^{k-1}=\frac{c_{i}^{k}}{k},\quad i=1,\ldots,s,\quad k=1,\ldots,s.$
Vectorized notation
We introduce notation to facilitate the description of the proposed vectorized implementation of the fixed-point iteration.
Let $b$ and $c$ denote the $s$-vectors $(b_{1},\ldots,b_{s})$ and $(c_{1},\ldots,c_{s})$, respectively. Similarly, for each $i=1,\ldots,s$, let $\mu_{i}$ and $\nu_{i}$ represent the $s$-vectors $(\mu_{1i},\ldots,\mu_{si})$ and $(\nu_{1i},\ldots,\nu_{si})$, respectively.
For an arbitrary $s$-vector $v=(v_{1},\ldots,v_{s})$ and a real number $\lambda\in\mathbb{R}$, we denote by $\lambda+v$ the $s$-vector $(\lambda+v_{1},\ldots,\lambda+v_{s})$. Given two arbitrary $s$-vectors $v=(v_{1},\ldots,v_{s})$ and $w=(w_{1},\ldots,w_{s})$, $v\circ w$ represents the componentwise product of two $s$-vectors defined by $(v_{1}w_{1},\ldots,v_{s}w_{s})$. We also use the notation $\mathrm{sum}(v)=v_{1}+\cdots+v_{s}$ and $\|v\|_{\infty}=\max(|v_{1}|,\ldots,|v_{s}|)$.
For each $j\in\{1,\ldots,D\}$, we denote by $Y_{n,i}^{j}$, $L_{n,i}^{j}$, and $f^{j}(t_{n-1}+hc_{i},Y_{n,i})$ the $j$th component of $Y_{n,i}\in\mathbb{R}^{D}$, $L_{n,i}\in\mathbb{R}^{D}$ and $f(t_{n-1}+hc_{i},Y_{n,i})\in\mathbb{R}^{D}$ respectively. For each $j\in\{1,\ldots,D\}$, we consider the $s$-vectors $Y_{n}^{j}=(Y_{n,1}^{j},\ldots,Y_{n,s}^{j})$ and $L_{n}^{j}=(L_{n,1}^{j},\ldots,L_{n,s}^{j})$.
Let $Y_{n}$ represent the one-dimensional array $(Y_{n}^{1},\ldots,Y_{n}^{D})$ of $s$-vectors $Y_{n}^{j}$. We define the function $F$ that given $t_{n-1}+hc$ and $Y_{n}$, returns the one-dimensional array $(F_{n}^{1},\ldots,F_{n}^{D})$ of $s$-vectors
$F_{n}^{j}=(f^{j}(t_{n-1}+hc_{1},Y_{n,1}),\ldots,f^{j}(t_{n-1}+hc_{s},Y_{n,s})).$
With that notation, (7) can be written as follows,
$\begin{split}F_{n}&=F(t_{n-1}+hc,Y_{n}),\\ L_{n}^{j}&=h\,(b\circ F_{n}^{j}),\quad j=1,\ldots,D,\\ Y_{n}^{j}&=y_{n-1}^{j}+\sum_{i=1}^{s}\mu_{i}L_{n,i}^{j},\quad j=1,\ldots,D,\end{split}$ (11)
while (8) can be rewritten as
$y_{n}^{j}=y_{n-1}^{j}+\mathrm{sum}(L_{n}^{j}),\quad j=1,\ldots,D.$ (12)
Vectorized implementation of fixed-point iteration
Following (9)–(10), we initialize the fixed-point iteration by approximating $Y_{n}^{j}$ for $j=1,\ldots,D$ as
$Y_{n}^{j,[0]}=y_{n-1}^{j}+\sum_{i=1}^{s}\nu_{i}L_{n-1,i}^{j}.$ (13)
Then, $Y_{n}^{[k]}=(Y_{n}^{1,[k]},\ldots,Y_{n}^{D,[k]})$ is computed for $k=1,2,\ldots$ as follows:
$\begin{split}F_{n}^{[k]}&=F(t_{n-1}+hc,Y_{n}^{[k-1]}),\\ L_{n}^{j,[k]}&=h\,(b\circ F_{n}^{j,[k]}),\quad j=1,\ldots,D,\\ Y_{n}^{j,[k]}&=y_{n-1}^{j}+\sum_{i=1}^{s}\mu_{i}L_{n,i}^{j,[k]},\quad j=1,\ldots,D.\end{split}$ (14)
Stopping criterion
In [12], it is shown that standard stopping criteria based on prescribed iteration error tolerances result in linear growth of energy error when applied to Hamiltonian systems. To address this issue, a new stopping criterion is proposed in [1]. We now present a simplified version of this criterion that is suitable for our vectorized implementation.
If $Y_{n}^{j,[k]}=Y_{n}^{j,[k-1]}$ for all $j\in\{1,\ldots,D\}$, the iteration should clearly be stopped; as $Y_{n}^{j,[k+1]}=Y_{n}^{j,[k]}$ would hold for all $j\in\{1,\ldots,D\}$. However, since this condition may not always be met, the stopping criterion must be supplemented with an additional condition to detect when successive approximations cease to improve: Define $\Delta_{n}^{j,[k]}=\|Y_{n}^{j,[k]}-Y_{n}^{j,[k-1]}\|_{\infty}$. The iteration will stop after the $k$th iteration if the following condition is met for all $j\in\{1,\ldots,D\}$:
$\Delta_{n}^{j,[k]}=0\quad\mbox{or}\quad\min\left(\Delta_{n}^{j,[1]},\ldots,\Delta_{n}^{j,[k-2]}\right)\leqslant\min\left(\Delta_{n}^{j,[k-1]},\Delta_{n}^{j,[k]}\right).$ (15)
Systems of second Order Differential Equations
We now consider systems of second-order differential equations of the form
$\frac{dq}{dt}=v,\quad\frac{dv}{dt}=g(q,t),\quad q(t_{0})=q_{0},\quad v(t_{0})=v_{0},$ (16)
where $q$ and $v$ are vectors with $d$ components each. Hamiltonian systems with a Hamiltonian function of the form
$H(q,p)=\frac{1}{2}\,p^{T}M^{-1}p+U(q)$ (17)
can be expressed in this form via the linear change of variables $p=Mv$.
System (16) is a specific instance of the general form (1), with $D=2d$, $y=(y^{1},\ldots,y^{2d})$, $q=(y^{1},\ldots,y^{d})$, $v=(y^{d+1},\ldots,y^{2d})$, and $f(t,y)=(v,g(t,q))$. It is known [18, 13], that a partitioned form of fixed-point iteration can generally reduce the number of iterations needed for convergence. Such partitioned fixed-point iteration also admits a vectorized implementation that we now describe.
We initialize the fixed-point iteration as
$Y_{n}^{j,[0]}=y_{n-1}^{j}+\sum_{i=1}^{s}\nu_{i}L_{n-1,i}^{j},\quad j=d+1,\ldots,2d.$ (18)
Then, for $k=1,2,\ldots$,
$\begin{split}L_{n}^{j,[k]}&=h\,(b\circ Y_{n}^{d+j,[k-1]}),\quad j=1,\ldots,d,\\ Y_{n}^{j,[k]}&=y_{n-1}^{j}+\sum_{i=1}^{s}\mu_{i}L_{n,i}^{j,[k]},\quad j=1,\ldots,d,\\ F_{n}^{[k]}&=F(t_{n-1}+hc,Y_{n}^{[k]}),\\ L_{n}^{j,[k]}&=h\,(b\circ F_{n}^{j,[k]}),\quad j=d+1,\ldots,2d,\\ Y_{n}^{j,[k]}&=y_{n-1}^{j}+\sum_{i=1}^{s}\mu_{i}L_{n,i}^{j,[k]},\quad j=d+1,\ldots,2d.\end{split}$ (19)
The process will stop after the $k$th iteration if (15) holds for all $j\in\{1,\ldots,d\}$.
3 Implementation aspects of IRKGL16-SIMD
3.1 SIMD-paralellization
Single Instruction, Multiple Data (SIMD) is a parallelization technique supported by modern CPU cores [17, 9]. SIMD enables the execution of vectorized instructions that apply a single operation simultaneously to multiple data elements, thereby accelerating performance in computationally intensive tasks. Modern CPUs contain specialized registers known as short vectors, typically 256 bits (holding four 64-bit double-precision floating-point numbers) or 512 bits (holding eight 64-bit double-precision floating-point numbers) in size.
In a standard SIMD operation, two input vectors are processed element-wise, applying the same operation to each pair of corresponding elements and producing an output vector. By operating on multiple elements in parallel, SIMD significantly improves performance.
3.2 Julia language and SIMD.jl
Our SIMD-vectorized implementation of IRKGL16 (IRKGL16-SIMD) is written in Julia language and relies on the SIMD.jl package. Julia is a high-level dynamic language that allows programmers to write clear, high-level, generic and abstract code resembling mathematical formulas, yet produces fast, low-level machine code that has traditionally only been generated by static languages [3, 4]. By default, Julia employs just-in-time (JIT) compilation, generating LLVM intermediate code that the LLVM compiler framework then translates into optimized machine code.
In Julia, there are several ways to explicitly apply SIMD vectorization, and the SIMD.jl package [19] provides a convenient solution for this purpose. The SIMD.jl library introduces the parameterized vector type Vec{s, T}, representing vectors of $s$ elements of type T. In our vectorized implementation of $s$-stage IRKGL methods (where $s=8$ in IRKGL16-SIMD), we use vectors of type Vec{s, Float64}, meaning $s$-vectors with elements of type ‘Float64‘ (64-bit double-precision floating-point numbers), or alternatively, vectors of type Vec{s, Float32}. The standard arithmetic and logical operations are designed to be applied element-wise in parallel, producing a SIMD vector as the result. SIMD.jl generates LLVM code that defines vectors of $s$ elements, which the LLVM compiler then translates into optimized SIMD instructions that utilize the SIMD registers available on the target platform.
3.3 VecArray and IRKGL16-SIMD
The vectorized form of fixed-point iteration (13)–(14) (resp., (18)–(19)) for first-order problems (1) (resp., second-order problems (16)) is well-suited for SIMD implementation. SIMD-vectorized implementation is particularly effective for IRKGL16, where $s=8$, as this choice maximizes efficient utilization of SIMD registers on modern hardware; in this case, all additions and multiplications in (13)–(14) and (18)–(19) can be implemented efficiently using SIMD operations with short vectors containing eight 64-bit double-precision floating-point numbers.
In addition, the evaluation of $F(t_{n-1}+ch,Y_{n}^{[k-1]})$ in (14) and (19) (which is equivalent to the evaluation of $f(t_{n-1}+c_{i}h,Y_{n,i}^{[k-1]})$ for $i=1,\ldots,s$) also benefits from SIMD parallelization: In IRKGL16-SIMD, parallelization remains transparent to the user, assuming the right-hand side of the system is defined in terms of arithmetic operations and elementary functions compatible with SIMD.jl. Such seamless SIMD parallelization is handled automatically with the help of SIMD.jl and an additional parametrized abstract type, VecArray{s,T}, that we have implemented to efficiently represent and handle abstract arrays of elements of type Vec{s,T} (where T is either Float64 or Float32). The $s$-vectors $b,c,\mu_{i},\nu_{i}$, $Y_{n}^{j}$, $F_{n}^{j}$, and $L_{n}^{j}$ considered in Subsection 2.4 are represented in IRKGL16-SIMD as vectors of type Vec{s,T}, while $Y_{n}=(Y_{n}^{1},\ldots,Y_{n}^{D})$, $F_{n}=(F_{n}^{1},\ldots,F_{n}^{D})$, and $L_{n}=(L_{n}^{1},\ldots,L_{n}^{D})$, and also $\mu=(\mu_{1},\ldots,\mu_{s})$ and $\nu=(\nu_{1},\ldots,\nu_{s})$, are represented as one-dimensional abstract arrays of type VecArray{s,T}.
Let us consider for instance the VecArray object Yn representing $Y_{n}$:
  • The syntax Yn[j] (or equivalently getindex(Yn,j)) can be used to get the vector of type Vec{s,T} representing $Y_{n}^{j}$.
  • The syntax Yn[j] = V (or equivalently setindex!(Yn,j,V)) can be used to update the $j$th vector of type Vec{s,T} representing $Y_{n}^{j}$. Here, we are assuming that V is a vector of type Vec{s,T}.
  • The data of Yn are internally stored as a 2-dimensional array of type Array{T,2} having $s$ rows and $D$ columns.
  • The $s$ components of Yn[j] are stored in the $j$th column of that 2-dimensional array.
Listing LABEL:code:ex1 shows the section of the IRKGL16-SIMD code where a single vectorized fixed-point iteration (14) is implemented.
1f_ODE!(Fn,Yn,parms,tn+h*c)
2
3for j in 1:D
4 Fnj=Fn[j]
5 Lnj=h*(b*Fnj)
6 dYnj=mu[1]*Lnj[1]
7 for i in 2:s
8 dYnj=dYnj+mu[i]*Lnj[i]
9 end
10 Yn[j]=yn[j]+dYnj
11end
The first line of this code requires some explanation, while the remaining lines in Listing LABEL:code:ex1 directly translate the last two lines of (14). The function f_ODE!(dy, y, parms, t) is a user-defined generic function that takes as inputs the array y, a vector parms of constant parameters, and the time t. It computes the $D$ components of the right-hand side $f(t,y)$ of the ODE system, storing the result in the pre-existing array dy. For illustration, Listing LABEL:code:ex2 provides an implementation of f_ODE!(dy, y, parms, t) for the following non-autonomous Hamiltonian ODE system:
$\begin{split}\frac{dq_{1}}{dt}&=p_{1},\quad\frac{dq_{2}}{dt}=p_{2},\\ \frac{dp_{1}}{dt}&=-q_{1}-2(\lambda+\xi\sin(t))q_{1}q_{2},\\ \frac{dp_{2}}{dt}&=-q_{2}-(\lambda+\xi\sin(t))(q_{1}^{2}-q_{2}^{2}).\end{split}$ (20)
Note that this system reduces to the classic Hénon-Heiles system when $\xi=0$. 1function f_ODE!(dy,y,parms,t) 2 3 # q1=y[1]; q2=y[2]; p1=y[3]; p2=y[4] 4 lambda=parms[1] 5 xi=parms[2] 6 aux=lambda+xi*sin(t) 7 dy[1]=y[3] 8 dy[2]=y[4] 9 dy[3]=-y[1]-2*aux*y[1]*y[2] 10 dy[4]=-y[2]-aux*(y[1]^2-y[2]^2) 11 12 return nothing 13end
The generic function f_ODE!(dy, y, parms, t) in Listing LABEL:code:ex2 supports a wide range of argument types. For example, it can be called with t as a scalar of type T = Float64 (or alternatively T=Float32 or T=BigFloat), and dy and y as standard Julia arrays containing elements of type T. Additionally, the function accepts other argument types, provided that the operations within the f_ODE! definition are valid for them. Specifically, the function call f_ODE!(Fn, Yn, parms, tn + h*c) in Listing LABEL:code:ex1, where Fn and Yn are abstract arrays of type VecArray{s, T} (with T = Float64 or T = Float32) and tn + h*c is of type Vec{s, T}, is equivalent to the first line of (14), representing the parallel evaluation of $f(t_{n-1}+c_{i}h,Y_{n,i}^{[k-1]})$ for $i=1,\ldots,s$.
This design enables parallel evaluations of the ODE system’s right-hand side in a way that is transparent to the user, provided that the right-hand side $f(t,y)$ of the ODE system is defined as a generic function f_ODE!(dy, y, parms, t) without branching in terms of arithmetic operations and elementary functions compatible with SIMD.jl. This requirement is met for all the problems considered in the numerical experiments in Section 5 below.
However, not all ODE systems (1) support a generic implementation of f_ODE!(dy, y, parms, t) that enables seamless $s$-fold parallel evaluations of the system’s right-hand side. For instance, this may not be feasible if the definition of $f(t,y)$ involves unavoidable branching, which disrupts the uniformity needed for SIMD processing. In such cases, however, IRKGL16-SIMD can still be applied if the user provides a specialized implementation of f_ODE!(dy, y, parms, t) for instances where dy and y are abstract arrays of type VecArray{s, T} (with T = Float64 or T = Float32) and t is of type Vec{s, T}.
In some cases, such as the $N$-body example considered in Section 5, it is more convenient for the user to organize the state variables in a higher-dimensional array rather than a one-dimensional array. To accommodate this, we have parameterized our abstract array type with an additional parameter dim, as VecArray{s, T, dim}. For one-dimensional arrays of vectors of type Vec{s, T}, we set dim = 2, with the underlying data stored in a two-dimensional standard array of type Array{T, 2}. More generally, VecArray{s, T, dim} refers to $(\texttt{dim}-1)$-dimensional abstract arrays of vectors of type Vec{s, T}, with the data stored in a dim-dimensional array of type Array{T, dim}. Our implementations of the getindex and setindex! functions work as expected in this more general case, allowing users to apply IRKGL16-SIMD with a generic implementation of the function f_ODE!(dy, y, parms, t) when the state variables are stored in a higher-dimensional array.
4 State-of-the-art explicit symplectic integrators
Splitting and composition techniques are powerful tools for constructing practical symplectic integrators in many areas of application. Although the resulting methods are typically tailored to the specific problem and lack the general applicability of approaches such as IRKGL schemes, they can be highly efficient when applicable. [18, 7, 11].
In the context of Hamiltonian systems that can be decomposed into two (or more) exactly solvable components, explicit symplectic integrators can be constructed via operator splitting techniques. A widely used approach is the second-order Strang splitting method [21], which yields a symmetric and symplectic scheme by composing the flows associated with the individual sub-Hamiltonians. Specifically, the Strang splitting scheme advances the solution over a time step $h$ via the composition
$\phi_{h}=\varphi^{A}_{h/2}\circ\varphi^{B}_{h}\circ\varphi^{A}_{h/2}.$ (21)
This method is second-order accurate, time-reversible, and symplectic. The resulting integrator is explicit when the Hamiltonian is written as $H=H_{A}+H_{B}$, and the exact flows $\varphi^{A}_{t}$ and $\varphi^{B}_{t}$ generated by $H_{A}$ and $H_{B}$, respectively, are available in closed form.
For instance, consider a separable Hamiltonian of the form (17), where $H_{A}(q,p)=\tfrac{1}{2}p^{T}M^{-1}p$ and $H_{B}(q,p)=U(q)$. In this case, the exact flows are given by $\varphi^{A}_{t}(q,p)=(q+tM^{-1}p,\,p)$ and $\varphi^{B}_{t}(q,p)=(q,\,p-t\nabla U(q))$.
The method can be extended to systems where the Hamiltonian splits into more than two exactly solvable components. For instance, if $H=H_{A}+H_{B}+H_{C}$ then a symmetric second-order integrator can be obtained by composing the flows of each part as
$\phi_{h}=\varphi^{A}_{h/2}\circ\varphi^{B}_{h/2}\circ\varphi^{C}_{h}\circ\varphi^{B}_{h/2}\circ\varphi^{A}_{h/2}.$ (22)
This generalized Strang-type splitting remains symplectic and time-symmetric, provided that each subflow $\varphi^{A}_{t}$, $\varphi^{B}_{t}$, and $\varphi^{C}_{t}$ is symplectic. Such multi-part splittings arise naturally in applications like charged particle dynamics, Lie-Poisson systems, and systems with coupled fast and slow Hamiltonian components [7, 14].
Extensions to higher-order accuracy can be obtained by symmetric compositions of $s$ basic Strang stages  [7, 11] of the form,
$\psi_{h}=\phi_{\gamma_{s}h}\circ\phi_{\gamma_{s-1}h}\circ\dots\circ\phi_{\gamma_{2}h}\circ\phi_{\gamma_{1}h}$ (23)
with $\gamma_{s+1-j}=\gamma_{j}$, $1\leqslant j\leqslant s$.
We consider the following symmetric compositions of time-symmetric second-order schemes of order $r=2,4,6,8,10$ recommended in [7]:
  • 1)
    Strang splitting: $r=2$th order with $s=1$ stage.
  • 2)
    SUZ90: $r=4$th order with $s=5$ stages proposed by Suzuki(1990) [22].
  • 3)
    SS05: a sixth-order scheme with $s=13$ stages, a eighth-order one with $s=21$ and a tenth-order one with $s=35$ proposed by Sofroniu and Spalleta(2005) [20].
Any composition of the form (23) with (21) can be rewritten as
$\varphi_{a_{s+1}h}^{A}\varphi_{b_{s}h}^{B}\circ\varphi_{a_{s}h}^{A}\circ\dots\circ\varphi_{b_{1}h}^{B}\circ\varphi_{a_{1}h}^{[A]}$ (24)
with coefficients
$a_{1}=\frac{\gamma_{1}}{2},\quad a_{j}=\frac{\gamma_{j-1}+\gamma_{j}}{2},\quad b_{j}=\gamma_{j}.$
Clearly, (24) cannot be written in the form (23) for arbitrary sequences of coefficients $a_{j},b_{j}$. Such more general family of methods are of particular interest in the case of Hamiltonians of the form (17), for which efficient high-order schemes have been constructed. They are often referred as splitting methods of Runge-Kutta-Nyström (RKN) type, as they can be rewritten as explicit symplectic RKN methods applied to the system of second order ODEs (16[18]. We consider the following splitting integrators of RKN type of order $r=6$ and $r=8$ recommended in [7]:
  • 1)
    $\mathrm{S_{ABA}\text{-}BM02}$: a sixth-order method with $s=14$ stages method proposed in Blanes and Moan(2002) [8].
  • 2)
    $\mathrm{S_{ABA}\text{-}BCE22}$: optimized eighth-order with $s=19$ stages proposed in Blanes, Casas and Escorihuela (2022) [6].
5 Numerical experiments
We compare the performance of IRKGL16-SIMD with several state-of-the-art explicit symplectic integrators for the high-precision integration of non-stiff Hamiltonian systems using constant step size. A range of representative examples is considered to assess both accuracy and efficiency.
Some of the explicit symplectic integrators used in the comparison are available through the Julia language’s DifferentialEquations.jl suite [16]. However, for consistency and fair benchmarking, we have implemented all the reference integrators described in the previous section. We verified that, for the problems considered, our implementations of the splitting methods are at least as efficient as—and often more so than—those available in DifferentialEquations.jl.
The source code for the integrators, along with all Jupyter notebooks used to generate the numerical results, is publicly available at: https://github.com/mikelehu/Implicit_symplectic_can_outperform_explicit_symplectic.
All numerical experiments were conducted on an 11th Gen Intel Core i7-11850H processor (2.5 GHz, 16 threads), a CPU equipped with 512-bit SIMD registers. We used the long-term support (LTS) stable release of Julia version 1.10.7 (November 26, 2024) for all tests. To optimize SIMD performance, Julia was launched from the terminal with bounds checking disabled using the flag --check-bounds=no.
5.1 Test problems
We have made comparisons with systems of first-order differential equations as well as with systems of second-order differential equations.
5.1.1 System of first-order differential equations
We consider a Hamiltonian system describing the motion of charged particles in the vicinity of a Schwarzschild black hole under the influence of an external magnetic field, as studied in [15]:
$\begin{split}H(r,\theta,p_{r},p_{\theta})&=\frac{1}{2}\left(1-\frac{2}{r}\right)p_{r}^{2}-\frac{1}{2}\left(1-\frac{2}{r}\right)^{-1}E^{2}+\frac{p_{\theta}^{2}}{2r^{2}}\\ &\quad+\frac{1}{2r^{2}\sin^{2}\theta}\left(L-\frac{\beta}{2}r^{2}\sin^{2}\theta\right)^{2}.\end{split}$ (25)
This Hamiltonian admits multiple splittings into analytically solvable components—typically into three, four, or five parts. Following the findings of [15], we adopt the three-part splitting
$H=H_{A}+H_{B}+H_{C},$
which was shown to offer the best numerical performance across both regular and chaotic trajectories. The individual components are defined as follows:
$H_{A}(r,\theta)$ $=\frac{1}{2r^{2}\sin^{2}\theta}\left(L-\frac{\beta}{2}r^{2}\sin^{2}\theta\right)^{2}-\frac{1}{2}\left(1-\frac{2}{r}\right)^{-1}E^{2},$
$H_{B}(r,p_{r},p_{\theta})$ $=\frac{1}{2}\left(p_{r}^{2}+\frac{p_{\theta}^{2}}{r^{2}}\right),$
$H_{C}(r,p_{r})$ $=-\frac{1}{r}p_{r}^{2}.$
Although [15] provides analytical expressions for the flows $\varphi_{t}^{A}$, $\varphi_{t}^{B}$, and $\varphi_{t}^{C}$, their direct implementation is computationally inefficient. To enable a fair and representative comparison with our implicit symplectic integrator (IRKGL16), we have improved the evaluation of these flows, significantly enhancing the performance of the explicit method. The derivation and implementation details of this optimized explicit integrator are presented in Appendix A.
As in [15], we set the parameters to $E=0.995$, $L=4.6$, and $\beta=8.9\times 10^{-4}$. The initial conditions are $\theta=\pi/2$ and $p_{r}=0$, with $r=11$ chosen so that the corresponding orbit is regular (i.e., lies on a closed curve). The initial value of $p_{\theta}>0$ is determined by imposing the energy condition $H=-\tfrac{1}{2}$.
5.1.2 Systems of second-order differential equations
We consider two second-order initial value problems of the form (16), corresponding to Hamiltonian systems of the type (17): the outer Solar System model [11] and the Hénon-Heiles problem [12].
6-body outer Solar System.
We examine a simplified model of the outer Solar System, which includes the Sun, the four outer planets (Jupiter, Saturn, Uranus, and Neptune), and Pluto, interacting via mutual Newtonian gravity. The system is described by an 18-degree-of-freedom Hamiltonian, with $q_{i},p_{i}\in\mathbb{R}^{3}$ for $i=1,\dots,6$, and Hamiltonian function
$H(q,p)=\sum_{i=1}^{N}\frac{\|p_{i}\|^{2}}{2m_{i}}-G\sum_{1\leq i<j\leq N}\frac{m_{i}m_{j}}{\|q_{i}-q_{j}\|},$ (26)
where $G$ is the gravitational constant, $m_{i}$ denotes the mass of body $i$, and $N=6$ is the total number of bodies considered.
Initial conditions are taken from the DE430 planetary ephemerides at Julian date 2440400.5 (June 28, 1969) [10], and adjusted so that the barycenter of the system is at rest. The integration is carried out over a time interval of $10^{7}$ days.
Hénon-Heiles system.
We also consider the classical Hénon-Heiles Hamiltonian:
$H(q,p)=\frac{1}{2}(p_{1}^{2}+p_{2}^{2})+\frac{1}{2}(q_{1}^{2}+q_{2}^{2})+q_{1}^{2}q_{2}-\frac{1}{3}q_{2}^{3}.$ (27)
This system is integrated from initial conditions corresponding to regular (non-chaotic) motion over a long time interval of length $2\pi\times 10^{4}$. Specifically, we choose $q_{1}(0)=0$, $q_{2}(0)=0.3$, $p_{2}(0)=0.2$, and determine $p_{1}(0)>0$ such that $H=1/12$ (see [11, 12]).
5.2 Results of Numerical Integrations
5.2.1 Systems of First-Order Differential Equations
Figure 1 presents work-precision diagrams for the Schwarzschild black hole problem. The left panel shows the maximum local error in the Hamiltonian,
$\Delta H^{\text{loc}}_{\max}=\max_{n}\left(\Delta H^{\text{loc}}_{n}\right),\quad\Delta H^{\text{loc}}_{n}=\left|\frac{H(y_{n})-H(y_{n-1})}{H(y_{n-1})}\right|,$ (28)
as a function of the total number of function evaluations. The right panel plots the same error against CPU time. In this plot, the minimal CPU time for the IRKGL16-SIMD method is highlighted with a red vertical dashed line.
We observe that even the sequential implementation of IRKGL16 outperforms the explicit symplectic integrators for the high precision integration of this problem. As for the SIMD implementation, for the same CPU time budget, it achieves significantly higher accuracy even for lower precision integrations.
Figure 1: Work-precision diagrams (maximum local error in the Hamiltonian) for the Schwarzschild black hole problem using the IRKGL16 method and symmetric composition schemes of orders $r=2,4,6,8$, and $10$. In the right plot, the sequential IRKGL16 implementation is denoted as IRKGL16-SEQ
Figure 2 shows the evolution of error over time for different integrators. Each method is integrated using a step size chosen such that the corresponding CPU time approximately matches the red vertical dashed line in the right panel of Figure 1. The left panel displays the global error in the Hamiltonian,
$\Delta H^{\text{glob}}_{\max}=\max_{n}\left(\Delta H^{\text{glob}}_{n}\right),\quad\Delta H^{\text{glob}}_{n}=\left|\frac{H(y_{n})-H(y_{0})}{H(y_{0})}\right|,$ (29)
while the right panel shows the relative error in the radial coordinate:
$\Delta r_{n}=\left|\frac{r_{n}-\tilde{r}_{n}}{\tilde{r}_{n}}\right|,$
where $\tilde{r}_{n}$ denotes the reference solution obtained using IRKGL16-SIMD with a comparatively small step size ($h=16$).
Figure 2: Left: global error in the Hamiltonian. Right: relative error in the radial coordinate $r$ for the Schwarzschild black hole problem. Results are shown for IRKGL16-SIMD and symmetric composition methods of orders $r=2,4,6,8,10$.
5.2.2 Systems of Second-Order Differential Equations
Figure 3 shows work-precision diagrams for the two second-order Hamiltonian problems. The top row corresponds to the 6-body outer Solar System model, while the bottom row pertains to the Hénon-Heiles system. The left panels display the maximum local Hamiltonian error (28) versus function evaluations, and the right panels show the same error versus CPU time. The red dashed line marks the optimal CPU time for the IRKGL16-SIMD method, in the sense that it is the minimal time required to achieve a local accuracy close to round-off error levels.
Figure 3: Work-precision diagrams (maximum local Hamiltonian error) for the 6-body outer Solar System and the Hénon-Heiles problem using IRKGL16, symmetric compositions (orders $r=2,4,6,8,10$), and symplectic RKN methods of orders 6 and 8.
Figures 46 show the evolution of errors, using simulations performed at optimal CPU time. The left panel in Figure 4 corresponds to the global Hamiltonian error of the 6-body outer Solar System, and the right panel to the Hénon-Heiles problem.
Figure 4: Time evolution of the global Hamiltonian error for the 6-body outer Solar System (left) and the Hénon-Heiles problem (right), using IRKGL16, symmetric compositions, and RKN methods.
Figure 5 depicts the position error of each planet in the outer Solar System over time, comparing IRKGL16 and SS05 ($r=10$). Figure 6 shows the error in both position and momentum for the Hénon-Heiles system across the different integrators.
Figure 5: Evolution of position error for each planet in the 6-body outer Solar System using IRKGL16 and SS05 ($r=10$).
Figure 6: Evolution of error in position (left) and momentum (right) for the Hénon-Heiles problem using IRKGL16, symmetric compositions, and RKN methods.
6 Conclusions
IRKGL methods are symplectic and, in non-stiff regimes, can be efficiently implemented using fixed-point iterations. These methods combine symplecticity, time-reversibility, and high-order of accuracy, and they are broadly applicable to general canonical Hamiltonian systems [18]. While optimized high-order explicit symplectic integrators are effective when the Hamiltonian can be decomposed into solvable components, they may be less efficient than IRKGL methods when high-precision integrations are required. The SIMD-based implementation of IRKGL16 significantly enhances performance, often outperforming specialized explicit symplectic methods even at moderate precision levels.
For non-stiff problems with Hamiltonians of the specific form (17), optimized high-order explicit symplectic integrators remain highly effective due to their simplicity and computational efficiency. Although our sequential implementation of IRKGL16 is less efficient than the best available explicit symplectic methods for standard double-precision computations, the SIMD-enhanced version can surpass these methods in performance for high-accuracy integrations, even when constrained to double-precision arithmetic.
[…]
Acknowledgments
All the authors have received funding by the Spanish State Research Agency through project PID2022-136585NB-C22, MCIN/AEI/10.13039/501100011033 and the European Union. They are also partially supported by the Department of Education of the Basque Government through the Consolidated Research Group MATHMODE (ITI456-22).
Appendix A Efficient implementation of the three-part splitting method for the Schwarzschild black hole example
The evaluation of the flows $\varphi_{t}^{A}$ and $\varphi_{t}^{B}$ becomes more efficient when performed in the variables $(x,y,p_{x},p_{y})$, which are related to the original coordinates by the canonical transformation
$x$ $=r\cos\theta,\quad y=r\sin\theta,$
$p_{x}$ $=\cos\theta\,p_{r}-\sin\theta\,\frac{p_{\theta}}{r},\quad p_{y}=\sin\theta\,p_{r}+\cos\theta\,\frac{p_{\theta}}{r}.$
In these new variables, the Hamiltonians $H_{A}$ and $H_{B}$ become
$\bar{H}_{A}(x,y)$ $=\frac{1}{2y^{2}}\left(L-\frac{\beta}{2}y^{2}\right)^{2}-\frac{1}{2}\left(1-\frac{2}{\sqrt{x^{2}+y^{2}}}\right)^{-1}E^{2},$
$\bar{H}_{B}(p_{x},p_{y})$ $=\frac{1}{2}(p_{x}^{2}+p_{y}^{2}).$
For the $H_{C}$ component, the equations of motion are simplified by introducing the variables $(\nu,\mu)$, defined through the non-canonical transformation
$\nu=r\,p_{r},\quad\mu=\frac{p_{r}^{2}}{r}.$
The corresponding equations are
$\frac{d\nu}{dt}=-3\mu,\quad\frac{d\mu}{dt}=0.$
The flows $\varphi_{t}^{A}$, $\varphi_{t}^{B}$, and $\varphi_{t}^{C}$—and hence the integrators (22) and (23)—can be efficiently evaluated in the variables $(x,y,p_{x},p_{y})$ as follows:
  • 1)
    Flow of $H_{A}$:
    $\varphi_{t}^{A}:(x,y,p_{x},p_{y})\mapsto(x,y,p_{x}^{*},p_{y}^{*}),$
    where
    $p_{x}^{*}$ $=p_{x}-t\,\frac{\partial\bar{H}_{A}}{\partial x}(x,y),$
    $p_{y}^{*}$ $=p_{y}-t\,\frac{\partial\bar{H}_{A}}{\partial y}(x,y).$
  • 2)
    Flow of $H_{B}$:
    $\varphi_{t}^{B}:(x,y,p_{x},p_{y})\mapsto(x^{*},y^{*},p_{x},p_{y}),$
    where
    $x^{*}$ $=x+t\,p_{x},\quad y^{*}=y+t\,p_{y}.$
  • 3)
    Flow of $H_{C}$:
    $\varphi_{t}^{C}:(x,y,p_{x},p_{y})\mapsto(x^{*},y^{*},p_{x}^{*},p_{y}^{*}),$
    where
    $r$ $=\sqrt{x^{2}+y^{2}},\quad c=\frac{x}{r},\quad s=\frac{y}{r},$
    $p_{\theta}$ $=-y\,p_{x}+x\,p_{y},\quad\nu=x\,p_{x}+y\,p_{y},\quad\mu=\frac{\nu^{2}}{r^{3}},$
    $\nu^{*}$ $=\nu-3t\,\mu,\quad p_{r}^{*}=(\mu\,\nu^{*})^{1/3},\quad r^{*}=\frac{\nu^{*}}{p_{r}^{*}},$
    $x^{*}$ $=r^{*}\,c,\quad y^{*}=r^{*}\,s,$
    $p_{x}^{*}$ $=c\,p_{r}^{*}-s\,\frac{p_{\theta}}{r^{*}},\quad p_{y}^{*}=s\,p_{r}^{*}+c\,\frac{p_{\theta}}{r^{*}}.$
The partial derivatives of $\bar{H}_{A}(x,y)$ can be computed efficiently using the intermediate variables:
$z$ $=y^{2},\quad r=\sqrt{x^{2}+z},\quad w=L-\frac{\beta}{2}z,\quad u=1-\frac{2}{r},$
$W$ $=-\frac{w}{z},\quad U=-\frac{E^{2}}{2u^{2}},\quad R=\frac{U}{r^{3}},\quad Z=R+\frac{(W-\beta)W}{2},$
$\frac{\partial\bar{H}_{A}}{\partial x}$ $=-2xR,\quad\frac{\partial\bar{H}_{A}}{\partial y}=-2yZ.$
Unlike the analytical expressions provided in [15] for the flows $\varphi_{t}^{A}$, $\varphi_{t}^{B}$, and $\varphi_{t}^{C}$, our implementation avoids the use of trigonometric and inverse trigonometric functions entirely, resulting in improved computational efficiency.

### v2

@@
 Abstract
-The main purpose of this work is to present a SIMD-vectorized implementation of the symplectic 16th-order 8-stage implicit Runge-Kutta integrator based on collocation with Gauss-Legendre nodes (IRKGL16-SIMD), and to show that it can outperform state-of-the-art symplectic explicit integrators for high-precision numerical integrations (in double-precision floating-point arithmetic) of non-stiff Hamiltonian ODE systems.
-Our IRKGL16-SIMD integrator leverages Single Instruction Multiple Data (SIMD) based parallelism (in a way that is transparent to the user) to significantly enhance the performance of the sequential IRKGL16 implementation.
-We present numerical experiments comparing IRKGL16-SIMD with state-of-the-art high-order explicit symplectic methods for the numerical integration of several Hamiltonian systems in double-precision floating-point arithmetic.
+Implicit Runge-Kutta schemes based on collocation with Gauss-Legendre nodes (IRKGL) possess attractive theoretical properties for long-term integration of Hamiltonian systems—they are symmetric, symplectic, and super-convergent—yet they have traditionally been considered less practical than explicit symplectic integrators.
+In this work, we challenge this conventional wisdom by introducing a stage-wise SIMD-vectorization approach that enables IRKGL schemes to outperform state-of-the-art explicit symplectic integrators for high-precision computations in double-precision floating-point arithmetic.
+Our approach reformulates the fixed-point iteration in terms of $s$-vectors (where $s$ is the number of stages) to explicitly exploit Single Instruction Multiple Data (SIMD) capabilities of modern processors, significantly reducing computational overhead and enabling parallel evaluation of the right-hand side function across all stages.
+Additionally, we present a reformulation of IRKGL schemes for second-order ODEs that ensures exact symplecticity at the level of double-precision floating-point arithmetic, extending our previous approach for first-order systems to this setting.
+We demonstrate these ideas through IRKGL16, a Julia implementation of the 16th-order 8-stage IRKGL scheme that performs vectorization seamlessly and transparently to the user.
+Numerical experiments on several Hamiltonian problems with separable structure confirm the effectiveness of our approach.
 keywords:
@@
 1 Introduction
-The family of $s$-stage implicit Runge-Kutta schemes based on collocation with Gauss-Legendre nodes (IRKGL) are known to be symmetric, symplectic and super-convergent of order $2s$, and hence are very well suited for high precision long-term numerical integration of Hamiltonian systems [18, 11].
-The computational cost per step of an implicit RK method is typically higher than that of an explicit RK method, as a system of algebraic equations must be solved at each step of implicit integrators.
-For non-stiff problems, fixed-point iterations are commonly employed to approximately solve such systems of algebraic equations at each step.
-A fixed-point implementation (based on the ideas presented in [1] and written in Julia language) of the $16$th order $8$-stage IRKGL scheme (IRKGL16) is included in the DifferentialEquations.jl suite [2, 16].
-Our code, IRKGL16, includes a standard time-step adaptivity mechanism.
-However, for long-term integration of Hamiltonian systems, the use of a constant time step is recommended [18, 11, 13].
-Therefore, in this work, we will focus on its application in constant step-size mode.
-While there are no explicit symplectic Runge-Kutta schemes, optimized high-order explicit symplectic integrators are available for Hamiltonian systems with suitable structure.
-These integrators, that can be derived with the techniques of splitting and composition, are applicable when the Hamiltonian function can be decomposed into the sum of two, three, or more exactly solvable Hamiltonians [5, 11, 13].
-It is widely acknowledged that high-order explicit symplectic methods are more efficient than IRKGL schemes for computations in double-precision floating-point arithmetic [11].
-In fact, to achieve similar accuracy over a given integration interval, our sequential implementation of IRKGL16 usually requires more CPU time than the most effective high order explicit symplectic integrators in double-precision computations.
-However, it is well established [11] that the efficiency of implicit Runge-Kutta methods can be significantly enhanced through various forms of parallelism.
-Single Instruction Multiple Data (SIMD), often referred to as SIMD-vectorization, is a type of parallelism that can be leveraged in most modern CPU cores, enabling vector computations at a cost comparable to scalar operations.
-The structure of the IRKGL16 integrator is particularly well-suited for exploiting SIMD parallelism with vectors containing eight (or alternatively, four or two) floating-point numbers.
-The primary objective of this work is to present a SIMD-vectorized implementation of IRKGL16 in constant step-size mode (referred to as IRKGL16-SIMD) and to demonstrate its superiority over state-of-the-art explicit symplectic integrators for high-precision numerical integrations in double-precision floating-point arithmetic for several relevant Hamiltonian problems with a separable structure.
-Our implementation makes use of the Julia package SIMD [19] to explicitly handle computations on vectors containing a fixed number of floating-point numbers.
-The algorithm IRKGL16-SIMD automatically performs vectorization, ensuring that the process remains transparent to the user.
-The integrator IRKGL16-SIMD is designed to numerically integrate non-stiff systems of (non-necessarily Hamiltonian) ODEs whose right-hand side is implemented as a generic Julia function using arithmetic operations and elementary functions supported by the SIMD package.
-The paper is organized as follows: Section 1 provides an introduction; Section 2 describes a vectorized fixed-point implementation of symplectic implicit Runge-Kutta (IRK) schemes; Section 3 presents implementation details of the IRKGL16-SIMD method; Section 4 reviews state-of-the-art symplectic explicit integrators; Section 5 reports numerical experiments; and Section 6 presents the conclusions.
+The family of $s$-stage implicit Runge-Kutta schemes based on collocation with Gauss-Legendre nodes (IRKGL) are known to be symmetric, symplectic and super-convergent of order $2s$, making them theoretically well-suited for high-precision long-term numerical integration of Hamiltonian systems [21, 13].
+However, despite these attractive theoretical properties, IRKGL schemes have traditionally been considered less practical than explicit symplectic integrators for Hamiltonian problems that admit a separable structure.
+The main reasons for this preference include: (i) the need for careful implementation to avoid linear drift of energy error [14, 1], (ii) the higher computational overhead inherent to implicit methods, and (iii) the typically larger number of right-hand side evaluations required to achieve comparable accuracy.
+For Hamiltonian systems with suitable structure—where the Hamiltonian function can be decomposed into the sum of two or more exactly solvable components—optimized high-order explicit symplectic integrators derived through splitting and composition techniques [8, 6, 13, 16] have thus become the methods of choice for high-precision integration.
+These methods avoid the complications associated with implicit schemes while maintaining the crucial symplectic property.
+In this work, we challenge this conventional wisdom by showing that IRKGL schemes can outperform state-of-the-art explicit symplectic integrators for high-precision computations in double-precision floating-point arithmetic.
+Our main contribution is a novel approach to formulating and implementing the fixed-point iterations of IRKGL schemes: we reformulate the iterations in terms of $s$-dimensional vectors (where $s$ is the number of stages) and design the algorithm to explicitly exploit Single Instruction Multiple Data (SIMD) capabilities of modern processors.
+This stage-wise SIMD-vectorization approach represents a departure from conventional strategies for reducing overhead in implicit integrators.
+Traditionally, SIMD-vectorization is achieved by structuring inner loops in a way that allows the compiler to automatically vectorize them.
+In contrast, our approach explicitly builds SIMD operations into the algorithm design itself.
+This has two crucial advantages: it significantly reduces the computational overhead of the implicit method, and it enables parallel evaluation of the right-hand side function across all $s$ stages simultaneously (provided the function satisfies certain requirements for vectorization).
+To ensure exact symplecticity at the level of double-precision floating-point arithmetic, we address the linear drift in energy error that can affect symplectic IRK schemes [14].
+For first-order ODEs, we adopt the approach presented in [1].
+Additionally, we generalize this approach to the implicit Runge-Kutta-Nyström (IRKN) formulation for second-order ODEs, which is the natural framework for many Hamiltonian systems arising from classical mechanics.
+We demonstrate these ideas through a concrete implementation, IRKGL16, which applies our stage-wise SIMD-vectorization strategy to the $16$th-order 88-stage IRKGL scheme (i.e., $s=8$, order $2s=16$).
+The implementation, written in Julia using the SIMD package [22], performs vectorization automatically and seamlessly, remaining transparent to the user.
+While IRKGL16 includes both constant step-size and adaptive step-size modes, in this work we focus on the constant step-size mode, as recommended for long-term symplectic integration of Hamiltonian systems [21, 13, 16].
+We have also implemented IRKGL methods with $s=2$, $s=4$, and $s=6$ stages, which are available in the IRKGL16 code, though our numerical experiments (see Subsection 5.2) demonstrate that the $8$-stage method achieves the best performance for high-precision computations.
+The code is designed for non-stiff systems of ODEs (not necessarily Hamiltonian).
+While it can handle arbitrary precision arithmetic, the vectorized implementation is available for single- and double-precision floating-point arithmetic.
+Maximum efficiency through full vectorization is achieved when the right-hand side function is implemented as a generic Julia function using arithmetic operations and elementary functions supported by the SIMD package.
+Through numerical experiments on several representative Hamiltonian problems with separable structure, we show that our stage-wise SIMD-vectorization approach enables IRKGL16 to outperform state-of-the-art explicit symplectic integrators for high-precision integration tasks in double-precision arithmetic.
+The paper is organized as follows: Section 2 describes the reformulation of IRKGL schemes for both first-order and second-order ODEs to ensure exact symplecticity and our stage-wise SIMD-vectorization strategy; Section 3 presents implementation details of IRKGL16; Section 4 reviews state-of-the-art high-order explicit symplectic integrators; Section 5 reports numerical experiments; and Section 6 presents the conclusions.
 2 Vectorized fixed-point implementation of symplectic IRK schemes
 2.1 Implicit RK methods
-We consider initial value problems for systems of ordinary differential equations (ODEs) of the form:
+In this work, we focus on the numerical integration of autonomous Hamiltonian systems of the form
+$\frac{dy}{dt}=J^{-1}\nabla H(y),\quad y\in\mathbb{R}^{2d},$ (1)
+where $H:\mathbb{R}^{2d}\to\mathbb{R}$ is the Hamiltonian function and $J$ is the standard symplectic matrix
+$J=\begin{pmatrix}0&I_{d}\\ -I_{d}&0\end{pmatrix}$.
+Runge-Kutta (RK) methods are widely used one-step integrators for numerically solving ODEs. These methods are specified by the coefficients in the Butcher tableau:
+[…]
+Although our focus is on Hamiltonian systems, RK methods are applicable to more general ODE problems of the more general form:
 $\frac{d}{dt}y=f(t,y),\quad y(t_{0})=y_{0},$ (2)
 where $f:\mathbb{R}^{D+1}\to\mathbb{R}^{D}$ is a sufficiently smooth function and $y_{0}\in\mathbb{R}^{D}$.
-Runge-Kutta (RK) methods are widely used one-step integrators for numerically solving ODEs. These methods are specified by the coefficients in the Butcher tableau:
-[…]
-The RK approximations $y_{n}\approx y(t_{n})$ to the solution $y(t)$ at times $t_{n}=t_{0}+nh$ for $n=1,2,3,\dots$ are given by:
+The RK approximations $y_{n}\approx y(t_{n})$ to the solution $y(t)$ of (2) at times $t=t_{n}=t_{0}+nh$ for $n=1,2,3,\dots$ are given by:
 $y_{n}=y_{n-1}+h\sum_{i=1}^{s}b_{i}f(t_{n-1}+c_{i}h,Y_{n,i}),$ (3)
@@
 Here, each $Y_{n,i}$ is an approximation of the state vector $y(t)$ at the intermediate time $t=t_{n-1}+hc_{i}$ computed within each time-step.
-The coefficients $a_{ij}$, $b_{i}$, and $c_{i}$ from the Butcher tableau determine the specific RK method and its properties.
-If the matrix $A=(a_{ij})$ is lower triangular the Runge-Kutta method is explicit.
+The vectors $b=(b_{1},\ldots,b_{s})$, $c=(c_{1},\ldots,c_{s})$, and the matrix $A=(a_{ij})$ from the Butcher tableau determine the specific RK method and its properties.
+If the matrix $A$ is lower triangular the Runge-Kutta method is explicit.
 Otherwise, it is an implicit Runge-Kutta (IRK) scheme, and the stage vectors $Y_{n,i}$ at each step are defined implicitly by the equation (4).
+The primary challenge in implementing implicit Runge-Kutta (IRK) methods is the efficient solution of the nonlinear system for the internal stages.
+For non-stiff problems, fixed-point iteration is generally recommended [21, 13].
 2.2 IRK methods of collocation type
-In a collocation method based on the nodes $c_{1},\ldots,c_{s}\in[0,1]$, the approximations $y_{n}\approx y(t_{n})$ to the solution $y(t)$ at times $t_{n}=t_{0}+nh$ are obtained as follows: for each $n=1,2,3,\dots$, consider the polynomial function $P_{n}:\mathbb{R}\to\mathbb{R}^{D}$ of degree $s$ uniquely determined by the following conditions,
+Among IRK methods, those based on collocation are particularly attractive.
+In a collocation method based on the pairwise distinct nodes $c_{1},\ldots,c_{s}\in[0,1]$, the approximations $y_{n}\approx y(t_{n})$ to the solution $y(t)$ at times $t_{n}=t_{0}+nh$ are obtained as follows: for each $n=1,2,3,\dots$, consider the polynomial function $P_{n}:\mathbb{R}\to\mathbb{R}^{D}$ of degree $s$ uniquely determined by the following conditions,
   • $P_{n}(t_{n-1})=y_{n-1}$,
@@
 then, set $y_{n}=P_{n}(t_{n})$.
-This is equivalent [11] to the application of the IRK method (2)–(3) with coefficients $b_{i},a_{ij}$ uniquely determined by the following conditions:
+This collocation procedure is equivalent [13] to the application of the IRK method (3)–(4) with coefficients $b_{i},a_{ij}$ uniquely determined by the following conditions:
 $\begin{split}\sum_{j=1}^{s}a_{ij}c_{j}^{k-1}&=\frac{c_{i}^{k}}{k},\quad i=1,\ldots,s,\quad k=1,2,\ldots,s,\\ \sum_{j=1}^{s}b_{j}c_{j}^{k-1}&=\frac{1}{k},\quad k=1,2,\ldots,s.\end{split}$ (5)
-In the specific case where $c_{1},\ldots,c_{s}\in(0,1)$ are the shifted Gauss-Legendre nodes–defined by $c_{i}=(1+x_{i})/2$, where $x_{1},\ldots,x_{s}$ are the zeroes of the Legendre polynomial of degree ss–the resulting IRK scheme of collocation type, which we refer to as the $s$-stage IRKGL scheme, achieves an order of accuracy $2s$ [11].
-2.3 Symplectic IRK methods with floating-point coefficients
-We consider an autonomous Hamiltonian system of the form
-$\frac{dy}{dt}=J^{-1}\nabla H(y),\quad y\in\mathbb{R}^{2d},$ (5)
-where $H:\mathbb{R}^{2d}\to\mathbb{R}$ is the Hamiltonian function and $J$ is the standard symplectic matrix
-$J=\begin{pmatrix}0&I_{d}\\ -I_{d}&0\end{pmatrix}$.
-The exact flow of this system, denoted by $\varphi_{H}^{t}$, is symplectic, meaning that it preserves the symplectic form: for all $t$, the Jacobian matrix $D\varphi_{H}^{t}(y)$ satisfies
-$\left(D\varphi_{H}^{t}(y)\right)^{T}JD\varphi_{H}^{t}(y)=J.$
-It is therefore desirable that numerical integrators used to approximate such flows also preserve this structure, in order to ensure long-time stability and accurate qualitative behavior of the solution.
-An implicit Runge-Kutta (IRK) method is symplectic if and only if [18] its coefficients satisfy the condition
-$b_{i}a_{ij}+b_{j}a_{ji}=b_{i}b_{j},\quad 1\leqslant i,j\leqslant s.$ (6)
-This condition ensures that the discrete flow defined by the IRK method is symplectic, and is satisfied in particular by the $s$-stage IRKGL schemes [18, 13].
-As shown in [12], if the coefficients $b_{i},a_{ij}$ are replaced by machine numbers $\tilde{b}_{i},\tilde{a}_{ij}$, then the resulting implicit RK scheme typically fails to satisfy the simplectic condition (6).
-To ensure that the symplectic property is exactly preserved with double-precision floating-point coefficients, we reformulate the IRKGL schemes as follows [1]:
-$Y_{n,i}$ $=y_{n-1}+\sum^{s}_{j=1}{\mu_{ij}\,L_{n,j},\quad L_{n,i}=hb_{i}f(t_{n-1}+hc_{i},Y_{n,i})},\quad i=1,\ldots,s,$ (7)
-$y_{n}$ $=y_{n-1}+\sum^{s}_{i=1}L_{n,i},$ (8)
-where $\mu_{ij}=a_{ij}/b_{j},\quad 1\leqslant i,j\leqslant s$.
-If the coefficients $b_{i}$, $c_{i}$, and $\mu_{ij}$ are replaced by the double-precision floating-point coefficients $\tilde{b}_{i}$, $\tilde{c}_{i}$, and $\tilde{\mu}_{ij}$, then the resulting scheme is still symplectic provided that $\tilde{\mu}_{ij}$ is set for $1\leqslant i<j\leqslant s$ as $1-\tilde{\mu}_{ji}$.
-2.4 Vectorized fixed-point iteration
-The primary challenge in implementing implicit Runge-Kutta (IRK) methods is the efficient solution of the nonlinear system for the internal stages.
-For non-stiff problems, fixed-point iteration is generally recommended.
-In collocation-type IRK methods, it is standard practice to initialize the fixed-point iteration with an estimate of the internal stages extrapolated from the previous time-step, that is,
+In the specific case where $c_{1},\ldots,c_{s}\in(0,1)$ are the shifted Gauss-Legendre nodes (defined by $c_{i}=(1+x_{i})/2$, where $x_{1},\ldots,x_{s}$ are the zeros of the Legendre polynomial of degree $s$) the resulting IRK scheme of collocation type, which we refer to as the $s$-stage IRKGL scheme, achieves an order of accuracy $2s$ [13].
+In collocation-type IRK methods, it is standard practice to initialize the fixed point iteration with an estimate of the internal stages extrapolated from the previous time-step, that is,
 $Y_{n,i}\approx P_{n-1}(t_{n-1}+hc_{i}),\quad i=1,\ldots,s.$ (6)
@@
 $\sum_{j=1}^{s}\nu_{ij}(c_{i}-1)^{k-1}=\frac{c_{i}^{k}}{k},\quad i=1,\ldots,s,\quad k=1,\ldots,s.$
+2.3 Application of IRK methods to second order differential equations
+Autonomous Hamiltonian systems with a Hamiltonian function of the form
+$H(q,p)=\frac{1}{2}\,p^{T}M^{-1}p+U(q)$ (8)
+can be expressed as system of second order ODEs of the form
+$\frac{d^{2}}{dt^{2}}q=M^{-1}\nabla U(q).$ (9)
+Any RK scheme can be applied to systems of second-order differential equations $\frac{d^{2}}{dt^{2}}q=g(t,q)$ by rewriting it as a system of first-order differential equations,
+$\frac{dq}{dt}=v,\quad\frac{dv}{dt}=g(q,t),$ (10)
+Obviously, the system (10) is a specific instance of the general form (2), with $D=2d$, $y=(y^{1},\ldots,y^{2d})$, $q=(y^{1},\ldots,y^{d})$, $v=(y^{d+1},\ldots,y^{2d})$, and $f(t,y)=(v,g(t,q))$.
+It is well known that the implicit equations of the IRK scheme can be simplified in the case of second order ODE systems of the form (10) by rewriting it as an implicit Runge-Kutta-Nystøm (IRKN) method.
+More precisely, the approximations $y_{n}=(q_{n},v_{n})\approx(q(t_{n}),v(t_{n}))$ of the solution $y(t)=(q(t),v(t))$ of the (10) supplemented with the initial conditions
+$q(t_{0})=q_{0},\quad v(t_{0})=v_{0}$ (11)
+can be computed as follows: for $n=1,2,\ldots$
+$\begin{split}q_{n}&=q_{n-1}+h\,v_{n-1}+h^{2}\sum_{i=1}^{s}\beta_{i}\,g(t_{n-1}+c_{i}h,Q_{n,i}),\\ v_{n}&=v_{n-1}+h\sum_{i=1}^{s}b_{i}g(t_{n-1}+c_{i}h,Q_{n,i}),\end{split}$ (12)
+where the vectors $Q_{n,i}$, $i=1,\ldots,s$, are implicitly defined by
+$Q_{n,i}=q_{n-1}+h\,c_{i}\,v_{n-1}+h^{2}\,\sum_{j=1}^{s}\alpha_{ij}\,g(t_{n-1}+c_{j}h,Q_{n,j}),\quad i=1,\ldots,s.$ (13)
+Here, $\beta_{i}=\sum_{k=1}^{s}b_{k}a_{ki}$, $c_{i}=\sum_{k=1}^{s}a_{ik}$, and $\alpha_{ij}=\sum_{k=1}^{s}a_{ik}a_{kj}$.
+This formulation gives rise to a more efficient fixed-point iteration, which, compared to the standard IRK formulation, approximately halves the required number of iterations for solving the implicit equations [21, 13].
+2.4 Symplectic IRK methods with floating-point coefficients
+Symplectic IRK methods
+The $t$-flow of the Hamiltonian system (1), denoted by $\varphi_{t}$, is symplectic, meaning that it preserves the symplectic form: for all $t$, the Jacobian matrix $D\varphi_{t}(y)$ satisfies
+$\left(D\varphi_{t}(y)\right)^{T}JD\varphi_{t}(y)=J.$
+It is therefore desirable that numerical integrators used to approximate such flows also preserve this structure, in order to ensure long-time stability and accurate qualitative behavior of the solution.
+An implicit Runge-Kutta (IRK) method is symplectic if and only if [21] its coefficients satisfy the condition
+$b_{i}a_{ij}+b_{j}a_{ji}=b_{i}b_{j},\quad 1\leqslant i,j\leqslant s.$ (14)
+This condition ensures that the discrete flow defined by the IRK method is symplectic, and is satisfied in particular by the $s$-stage IRKGL schemes [21, 16].
+Symplectic IRKN methods
+An IRK method rewritten as an IRKN (12)–(13) is symplectic if its coefficients satisfy the conditions [25, 21]
+$\begin{split}\beta_{i}&=b_{i}\,(1-c_{i}),\quad 1\leqslant i\leqslant s,\\ b_{i}\,(\beta_{j}-\alpha_{ij})&=b_{j}\,(\beta_{i}-\alpha_{ji}),\quad 1\leqslant i,j\leqslant s.\end{split}$ (15)
+Reformulation for floating-point arithmetic
+As shown in [14], if the coefficients $b_{i},a_{ij}$ of the IRK schemes (or $b_{i},\beta_{i},c_{i},\alpha_{ij}$ of the IRKN formulation) are replaced by their machine number representatives, then the resulting implicit RK scheme typically fails to satisfy the symplecticity condition (14).
+To ensure that the symplectic property is exactly preserved with double-precision floating-point coefficients, for first-order systems, we adopt the reformulation from [1]:
+$Y_{n,i}$ $=y_{n-1}+\sum^{s}_{j=1}{\mu_{ij}\,L_{n,j},\quad L_{n,i}=hb_{i}f(t_{n-1}+hc_{i},Y_{n,i})},\quad i=1,\ldots,s,$ (16)
+$y_{n}$ $=y_{n-1}+\sum^{s}_{i=1}L_{n,i},$ (17)
+where $b_{i}$ and $c_{i}$ are the machine-number representatives of the original coefficients of the IRK scheme, and
+$\mu_{ij}$ $=\mathrm{fl}(a_{ij}/b_{j}),\quad 1\leqslant j\leqslant i\leqslant s,$
+$\mu_{ij}$ $=1-\mu_{ji},\quad 1\leqslant i<j\leqslant s.$
+Here, $\mathrm{fl}(x)$ represents the machine-number representative of the real number $x$.
+For second-order systems, we generalize this approach as follows: We reformulate the IRKN form (12)–(13) of IRKGL schemes in terms of the coefficients $b_{i}$, $c_{i}$, and $\eta_{ij}=\alpha_{ij}/b_{j}$, so that the second condition in (15) can be replaced by
+$\eta_{ij}+c_{j}=\eta_{ji}+c_{i}.$ (18)
+Now, we define one step of the method as follows:
+$G_{n,i}$ $=g(t_{n-1}+hc_{i},Q_{n,i}),\quad R_{n,i}=h\,b_{i}\,G_{n,i},\quad i=1,\ldots,s,$ (19)
+$Q_{n,i}$ $=q_{n-1}+h\,c_{i}\,v_{n-1}+h\sum^{s}_{j=1}\eta_{ij}\,R_{n,j},\quad i=1,\ldots,s,$ (20)
+$v_{n}$ $=v_{n-1}+\sum^{s}_{i=1}R_{n,i},\quad q_{n}=q_{n-1}+h\,v_{n}-h\sum^{s}_{i=1}c_{i}R_{n,i}$ (21)
+where the coefficients $b_{i},c_{i}$ are machine-number representatives of the original coefficients of the $s$-stage IRKGL scheme, and
+$\eta_{ij}$ $=\mathrm{fl}(\alpha_{ij}/b_{j}),\quad 1\leqslant j\leqslant i\leqslant s,$
+$\eta_{ij}$ $=c_{i}-c_{j}-\eta_{ji},\quad 1\leqslant i<j\leqslant s.$
+This guarantees that the symplecticity conditions (18) holds.
+We have verified numerically that the coefficients $\eta_{ij}$ computed by this procedure are exactly representable as machine numbers in double-precision arithmetic for $s\leqslant 16$.
+2.5 Stage-wise vectorization of fixed-point iteration
+In order to enable SIMD implementation, we reformulate the method using $s$-vectors that group values across all stages.
 Vectorized notation
-We introduce notation to facilitate the description of the proposed vectorized implementation of the fixed-point iteration.
-Let $b$ and $c$ denote the $s$-vectors $(b_{1},\ldots,b_{s})$ and $(c_{1},\ldots,c_{s})$, respectively.
-Similarly, for each $i=1,\ldots,s$, let $\mu_{i}$ and $\nu_{i}$ represent the $s$-vectors $(\mu_{1i},\ldots,\mu_{si})$ and $(\nu_{1i},\ldots,\nu_{si})$, respectively.
-For an arbitrary $s$-vector $v=(v_{1},\ldots,v_{s})$ and a real number $\lambda\in\mathbb{R}$, we denote by $\lambda+v$ the $s$-vector $(\lambda+v_{1},\ldots,\lambda+v_{s})$.
-Given two arbitrary $s$-vectors $v=(v_{1},\ldots,v_{s})$ and $w=(w_{1},\ldots,w_{s})$, $v\circ w$ represents the componentwise product of two $s$-vectors defined by $(v_{1}w_{1},\ldots,v_{s}w_{s})$.
-We also use the notation $\mathrm{sum}(v)=v_{1}+\cdots+v_{s}$ and $\|v\|_{\infty}=\max(|v_{1}|,\ldots,|v_{s}|)$.
-For each $j\in\{1,\ldots,D\}$, we denote by $Y_{n,i}^{j}$, $L_{n,i}^{j}$, and $f^{j}(t_{n-1}+hc_{i},Y_{n,i})$ the $j$th component of $Y_{n,i}\in\mathbb{R}^{D}$, $L_{n,i}\in\mathbb{R}^{D}$ and $f(t_{n-1}+hc_{i},Y_{n,i})\in\mathbb{R}^{D}$ respectively.
-For each $j\in\{1,\ldots,D\}$, we consider the $s$-vectors $Y_{n}^{j}=(Y_{n,1}^{j},\ldots,Y_{n,s}^{j})$ and $L_{n}^{j}=(L_{n,1}^{j},\ldots,L_{n,s}^{j})$.
-Let $Y_{n}$ represent the one-dimensional array $(Y_{n}^{1},\ldots,Y_{n}^{D})$ of $s$-vectors $Y_{n}^{j}$.
-We define the function $F$ that given $t_{n-1}+hc$ and $Y_{n}$, returns the one-dimensional array $(F_{n}^{1},\ldots,F_{n}^{D})$ of $s$-vectors
-$F_{n}^{j}=(f^{j}(t_{n-1}+hc_{1},Y_{n,1}),\ldots,f^{j}(t_{n-1}+hc_{s},Y_{n,s})).$
+We introduce notation to facilitate the description of the proposed stage-wise vectorized implementation of the fixed-point iteration.
+Let $\boldsymbol{b}$ and $\boldsymbol{c}$ denote the $s$-vectors $(b_{1},\ldots,b_{s})$ and $(c_{1},\ldots,c_{s})$, respectively.
+Similarly, for each $i=1,\ldots,s$, let $\boldsymbol{\mu_{i}}$, $\boldsymbol{\nu_{i}}$, and $\boldsymbol{\eta_{i}}$ represent the $s$-vectors $(\mu_{1i},\ldots,\mu_{si})$, $(\nu_{1i},\ldots,\nu_{si})$, and $(\eta_{1i},\ldots,\eta_{si})$, respectively.
+For an arbitrary $s$-vector $\boldsymbol{v}=(v_{1},\ldots,v_{s})$ and a real number $\lambda\in\mathbb{R}$, we denote by $\lambda+\boldsymbol{v}$ the $s$-vector $(\lambda+v_{1},\ldots,\lambda+v_{s})$.
+Given two arbitrary $s$-vectors $\boldsymbol{v}=(v_{1},\ldots,v_{s})$ and $\boldsymbol{w}=(w_{1},\ldots,w_{s})$, $\boldsymbol{v}*\boldsymbol{w}$ represents the componentwise product of two $s$-vectors defined by $(v_{1}w_{1},\ldots,v_{s}w_{s})$.
+We also use the notation $\mathrm{sum}(\boldsymbol{v})=v_{1}+\cdots+v_{s}$, and $\|\boldsymbol{v}\|_{\infty}=\max(|v_{1}|,\ldots,|v_{s}|)$.
+To distinguish $s$-vectors from collections of $D$ state variables, we denote the laters as
+$y=\left[\begin{matrix}y^{1}\\ \vdots\\ y^{D}\end{matrix}\right]$
+(or alternatively as $y=[y^{1},\ldots,y^{D}]$) and refer to them as arrays of state variables.
+For each $j\in\{1,\ldots,D\}$, we denote by $Y_{n,i}^{j}$, $L_{n,i}^{j}$, and $f^{j}(t_{n-1}+hc_{i},Y_{n,i})$ the $j$th component of the arrays $Y_{n,i}\in\mathbb{R}^{D}$, $L_{n,i}\in\mathbb{R}^{D}$ and $f(t_{n-1}+hc_{i},Y_{n,i})\in\mathbb{R}^{D}$ respectively.
+For each $j\in\{1,\ldots,D\}$, we consider the $s$-vectors $\boldsymbol{Y_{n}^{j}}=(Y_{n,1}^{j},\ldots,Y_{n,s}^{j})$ and $\boldsymbol{L_{n}^{j}}=(L_{n,1}^{j},\ldots,L_{n,s}^{j})$.
+Let $\boldsymbol{Y_{n}}$ and $\boldsymbol{F_{n}}$ represent the arrays of $s$-vectors $\boldsymbol{Y_{n}^{j}}$ and $\boldsymbol{F_{n}^{j}}$
+$\boldsymbol{Y_{n}}=\left[\begin{matrix}\boldsymbol{Y_{n}^{1}}\\ \vdots\\ \boldsymbol{Y_{n}^{D}}\end{matrix}\right],\quad\boldsymbol{F_{n}}=\left[\begin{matrix}\boldsymbol{F_{n}^{1}}\\ \vdots\\ \boldsymbol{F_{n}^{D}}\end{matrix}\right].$
+We define the vectorized function $\boldsymbol{f}$ that given $t_{n-1}+h\boldsymbol{c}$ and $\boldsymbol{Y_{n}}$, returns the array $\boldsymbol{F_{n}}$ of $s$-vectors
+$\boldsymbol{F_{n}^{j}}=(f^{j}(t_{n-1}+hc_{1},Y_{n,1}),\ldots,f^{j}(t_{n-1}+hc_{s},Y_{n,s})),\quad j=1,\ldots,D.$
+That function $\boldsymbol{f}$ represents the parallel evaluation of $f(t_{n-1}+hc_{1},Y_{n,1}),\ldots,f(t_{n-1}+hc_{s},Y_{n,s})$.
+For instance, if $D=2$, $s=2$, and $f(t,[y^{1},y^{2}])=[2\,y^{1}*y^{2},t*(y^{1}+y^{2})]$, then $\boldsymbol{f}(t+h\boldsymbol{c},\boldsymbol{Y_{n}})$ returns the array
+$\left[\begin{matrix}2\,\boldsymbol{Y_{n}}^{1}*\boldsymbol{Y_{n}}^{2}\\ (t+h\boldsymbol{c})*(\boldsymbol{Y_{n}}^{1}+\boldsymbol{Y_{n}}^{2})\end{matrix}\right]=\left[\begin{matrix}(2\,Y_{n,1}^{1}*Y_{n,1}^{2},Y_{n,2}^{1}*Y_{n,2}^{2})\\ ((t+hc_{1})*(Y_{n,1}^{1}+Y_{n,1}^{2}),(t+hc_{2})*(Y_{n,2}^{1}+Y_{n,2}^{2}))\end{matrix}\right]$
+Note how operations on $s$-vectors are performed component-wise, enabling parallel evaluation across all stages.
 With that notation, (16) can be written as follows,
-$\begin{split}F_{n}&=F(t_{n-1}+hc,Y_{n}),\\ L_{n}^{j}&=h\,(b\circ F_{n}^{j}),\quad j=1,\ldots,D,\\ Y_{n}^{j}&=y_{n-1}^{j}+\sum_{i=1}^{s}\mu_{i}L_{n,i}^{j},\quad j=1,\ldots,D,\end{split}$ (11)
+$\begin{split}\boldsymbol{F_{n}}&=\boldsymbol{f}(t_{n-1}+h\,\boldsymbol{c},\boldsymbol{Y_{n}}),\\ \boldsymbol{L_{n}^{j}}&=h\,(\boldsymbol{b}*\boldsymbol{F_{n}^{j}}),\quad j=1,\ldots,D,\\ \boldsymbol{Y_{n}^{j}}&=y_{n-1}^{j}+\sum_{i=1}^{s}\boldsymbol{\mu_{i}}\,L_{n,i}^{j},\quad j=1,\ldots,D,\end{split}$ (22)
 while (17) can be rewritten as
-$y_{n}^{j}=y_{n-1}^{j}+\mathrm{sum}(L_{n}^{j}),\quad j=1,\ldots,D.$ (12)
+$y_{n}^{j}=y_{n-1}^{j}+\mathrm{sum}(\boldsymbol{L_{n}^{j}}),\quad j=1,\ldots,D.$ (23)
 Vectorized implementation of fixed-point iteration
-Following (9)–(10), we initialize the fixed-point iteration by approximating $Y_{n}^{j}$ for $j=1,\ldots,D$ as
-$Y_{n}^{j,[0]}=y_{n-1}^{j}+\sum_{i=1}^{s}\nu_{i}L_{n-1,i}^{j}.$ (13)
-Then, $Y_{n}^{[k]}=(Y_{n}^{1,[k]},\ldots,Y_{n}^{D,[k]})$ is computed for $k=1,2,\ldots$ as follows:
-$\begin{split}F_{n}^{[k]}&=F(t_{n-1}+hc,Y_{n}^{[k-1]}),\\ L_{n}^{j,[k]}&=h\,(b\circ F_{n}^{j,[k]}),\quad j=1,\ldots,D,\\ Y_{n}^{j,[k]}&=y_{n-1}^{j}+\sum_{i=1}^{s}\mu_{i}L_{n,i}^{j,[k]},\quad j=1,\ldots,D.\end{split}$ (14)
+Following (6)–(7), we initialize the fixed-point iteration by approximating $\boldsymbol{Y_{n}^{j}}$ for $j=1,\ldots,D$ as
+$\boldsymbol{Y_{n}^{j,[0]}}=y_{n-1}^{j}+\sum_{i=1}^{s}\boldsymbol{\nu_{i}}\,L_{n-1,i}^{j}.$ (24)
+Then, $\boldsymbol{Y_{n}^{[k]}}=(Y_{n}^{1,[k]},\ldots,Y_{n}^{D,[k]})$ is computed for $k=1,2,\ldots$ as follows:
+$\begin{split}\boldsymbol{F_{n}^{[k]}}&=\boldsymbol{f}(t_{n-1}+h\,\boldsymbol{c},\boldsymbol{Y_{n}^{[k-1]}}),\\ \boldsymbol{L_{n}^{j,[k]}}&=h\,(\boldsymbol{b}*\boldsymbol{F_{n}^{j,[k]}}),\quad j=1,\ldots,D,\\ \boldsymbol{Y_{n}^{j,[k]}}&=y_{n-1}^{j}+\sum_{i=1}^{s}\boldsymbol{\mu_{i}}\,L_{n,i}^{j,[k]},\quad j=1,\ldots,D.\end{split}$ (25)
+This formulation naturally enables SIMD implementation: the $s$-vectors can be stored in SIMD registers, and operations like $\boldsymbol{b}*\boldsymbol{F_{n}^{j}}$ correspond to single SIMD instructions.
+As for the alternative IRKN form (19)–(21) of IRKGL methods for second order ODE systems, we initialize the fixed-point iteration as
+$\begin{split}\boldsymbol{Y_{n}^{j,[0]}}&=y_{n-1}^{j}+\sum_{i=1}^{s}\boldsymbol{\nu_{i}}\,L_{n-1,i}^{j},\quad j=d+1,\ldots,2d,\\ \boldsymbol{L_{n}^{j,[0]}}&=h\,(\boldsymbol{b}*\boldsymbol{Y_{n}^{d+j,[0]}}),\quad j=1,\ldots,d,\\ \boldsymbol{Y_{n}^{j,[0]}}&=y_{n-1}^{j}+\sum_{i=1}^{s}\boldsymbol{\mu_{i}}L_{n,i}^{j,[0]},\quad j=1,\ldots,d.\end{split}$ (26)
+Then, for $k=1,2,\ldots$,
+$\begin{split}\boldsymbol{F_{n}^{[k]}}&=\boldsymbol{f}(t_{n-1}+h\,\boldsymbol{c},Y_{n}^{[k-1]}),\\ \boldsymbol{L_{n}^{j,[k]}}&=h\,(\boldsymbol{b}*\boldsymbol{F_{n}^{j,[k]}}),\quad j=d+1,\ldots,2d,\\ \boldsymbol{Y_{n}^{j,[k]}}&=y_{n-1}^{j}+h\,\left(\boldsymbol{c}\,y_{n-1}^{d+j}+\sum_{i=1}^{s}\boldsymbol{\eta_{i}}\,L_{n,i}^{d+j,[k]}\right),\quad j=1,\ldots,d,\end{split}$ (27)
 Stopping criterion
@@
 We now present a simplified version of this criterion that is suitable for our vectorized implementation.
-If $Y_{n}^{j,[k]}=Y_{n}^{j,[k-1]}$ for all $j\in\{1,\ldots,D\}$, the iteration should clearly be stopped; as $Y_{n}^{j,[k+1]}=Y_{n}^{j,[k]}$ would hold for all $j\in\{1,\ldots,D\}$.
-However, since this condition may not always be met, the stopping criterion must be supplemented with an additional condition to detect when successive approximations cease to improve: Define $\Delta_{n}^{j,[k]}=\|Y_{n}^{j,[k]}-Y_{n}^{j,[k-1]}\|_{\infty}$.
+If $\boldsymbol{Y_{n}^{j,[k]}}=\boldsymbol{Y_{n}^{j,[k-1]}}$ for all $j\in\{1,\ldots,D\}$, the iteration should clearly be stopped; as $\boldsymbol{Y_{n}^{j,[k+1]}}=\boldsymbol{Y_{n}^{j,[k]}}$ would hold for all $j\in\{1,\ldots,D\}$.
+However, since this condition may not always be met, the stopping criterion must be supplemented with an additional condition to detect when successive approximations cease to improve: Define
+$\Delta_{n}^{j,[k]}=\|\boldsymbol{Y_{n}^{j,[k]}}-\boldsymbol{Y_{n}^{j,[k-1]}}\|_{\infty}.$ (28)
 The iteration will stop after the $k$th iteration if the following condition is met for all $j\in\{1,\ldots,D\}$:
 $\Delta_{n}^{j,[k]}=0\quad\mbox{or}\quad\min\left(\Delta_{n}^{j,[1]},\ldots,\Delta_{n}^{j,[k-2]}\right)\leqslant\min\left(\Delta_{n}^{j,[k-1]},\Delta_{n}^{j,[k]}\right).$ (29)
-Systems of second Order Differential Equations
-We now consider systems of second-order differential equations of the form
-$\frac{dq}{dt}=v,\quad\frac{dv}{dt}=g(q,t),\quad q(t_{0})=q_{0},\quad v(t_{0})=v_{0},$ (16)
-where $q$ and $v$ are vectors with $d$ components each. Hamiltonian systems with a Hamiltonian function of the form
-$H(q,p)=\frac{1}{2}\,p^{T}M^{-1}p+U(q)$ (17)
-can be expressed in this form via the linear change of variables $p=Mv$.
-System (16) is a specific instance of the general form (1), with $D=2d$, $y=(y^{1},\ldots,y^{2d})$, $q=(y^{1},\ldots,y^{d})$, $v=(y^{d+1},\ldots,y^{2d})$, and $f(t,y)=(v,g(t,q))$.
-It is known [18, 13], that a partitioned form of fixed-point iteration can generally reduce the number of iterations needed for convergence.
-Such partitioned fixed-point iteration also admits a vectorized implementation that we now describe.
-We initialize the fixed-point iteration as
-$Y_{n}^{j,[0]}=y_{n-1}^{j}+\sum_{i=1}^{s}\nu_{i}L_{n-1,i}^{j},\quad j=d+1,\ldots,2d.$ (18)
-Then, for $k=1,2,\ldots$,
-$\begin{split}L_{n}^{j,[k]}&=h\,(b\circ Y_{n}^{d+j,[k-1]}),\quad j=1,\ldots,d,\\ Y_{n}^{j,[k]}&=y_{n-1}^{j}+\sum_{i=1}^{s}\mu_{i}L_{n,i}^{j,[k]},\quad j=1,\ldots,d,\\ F_{n}^{[k]}&=F(t_{n-1}+hc,Y_{n}^{[k]}),\\ L_{n}^{j,[k]}&=h\,(b\circ F_{n}^{j,[k]}),\quad j=d+1,\ldots,2d,\\ Y_{n}^{j,[k]}&=y_{n-1}^{j}+\sum_{i=1}^{s}\mu_{i}L_{n,i}^{j,[k]},\quad j=d+1,\ldots,2d.\end{split}$ (19)
-The process will stop after the $k$th iteration if (15) holds for all $j\in\{1,\ldots,d\}$.
-3 Implementation aspects of IRKGL16-SIMD
+For second order ODE systems, condition (29) is checked for all $j\in\{1,\ldots,d\}$.
+3 Implementation aspects of IRKGL16
 3.1 SIMD-paralellization
@@
 By operating on multiple elements in parallel, SIMD significantly improves performance.
-3.2 Julia language and SIMD.jl
-Our SIMD-vectorized implementation of IRKGL16 (IRKGL16-SIMD) is written in Julia language and relies on the SIMD.jl package.
+3.2 Julia language and the package SIMD.jl
+Our SIMD-vectorized implementation of IRKGL16 is written in Julia language and relies on the package SIMD.jl [22].
 Julia is a high-level dynamic language that allows programmers to write clear, high-level, generic and abstract code resembling mathematical formulas, yet produces fast, low-level machine code that has traditionally only been generated by static languages [4, 5].
 By default, Julia employs just-in-time (JIT) compilation, generating LLVM intermediate code that the LLVM compiler framework then translates into optimized machine code.
-In Julia, there are several ways to explicitly apply SIMD vectorization, and the SIMD.jl package [19] provides a convenient solution for this purpose.
-The SIMD.jl library introduces the parameterized vector type Vec{s, T}, representing vectors of $s$ elements of type T.
-In our vectorized implementation of $s$-stage IRKGL methods (where $s=8$ in IRKGL16-SIMD), we use vectors of type Vec{s, Float64}, meaning $s$-vectors with elements of type ‘Float64‘ (64-bit double-precision floating-point numbers), or alternatively, vectors of type Vec{s, Float32}.
+In Julia, there are several ways to explicitly apply SIMD vectorization, and the package SIMD.jl provides a convenient solution for this purpose.
+That package introduces the parameterized vector type Vec{s, T}, representing vectors of $s$ elements of type T.
+In our vectorized implementation of $s$-stage IRKGL methods (where $s=8$ in IRKGL16), we use vectors of type Vec{s, Float64}, meaning $s$-vectors with elements of type Float64 (64-bit double-precision floating-point numbers), or alternatively, vectors of type Vec{s, Float32}.
 The standard arithmetic and logical operations are designed to be applied element-wise in parallel, producing a SIMD vector as the result.
-SIMD.jl generates LLVM code that defines vectors of $s$ elements, which the LLVM compiler then translates into optimized SIMD instructions that utilize the SIMD registers available on the target platform.
-3.3 VecArray and IRKGL16-SIMD
+The package SIMD.jl generates LLVM code that defines vectors of $s$ elements, which the LLVM compiler then translates into optimized instructions that utilize the SIMD registers available on the target platform, while also enabling other low-level optimizations (e.g., instruction-level parallelism) afforded by the Vec{s,T} abstraction.
+3.3 VecArray and IRKGL16
 The vectorized form of fixed-point iteration (24)–(25) (resp., (26)–(27)) for first-order problems (2) (resp., second-order problems (10)) is well-suited for SIMD implementation.
-SIMD-vectorized implementation is particularly effective for IRKGL16, where $s=8$, as this choice maximizes efficient utilization of SIMD registers on modern hardware; in this case, all additions and multiplications in (13)–(14) and (18)–(19) can be implemented efficiently using SIMD operations with short vectors containing eight 64-bit double-precision floating-point numbers.
-In addition, the evaluation of $F(t_{n-1}+ch,Y_{n}^{[k-1]})$ in (14) and (19) (which is equivalent to the evaluation of $f(t_{n-1}+c_{i}h,Y_{n,i}^{[k-1]})$ for $i=1,\ldots,s$) also benefits from SIMD parallelization: In IRKGL16-SIMD, parallelization remains transparent to the user, assuming the right-hand side of the system is defined in terms of arithmetic operations and elementary functions compatible with SIMD.jl.
+SIMD-vectorized implementation is particularly effective for IRKGL16, where $s=8$, as this choice maximizes efficient utilization of SIMD registers on modern hardware; in this case, all additions and multiplications in (24)–(25) and (26)–(27), as well as the computations in (28), can be implemented efficiently using SIMD operations with short vectors containing eight 64-bit double-precision floating-point numbers.
+In addition, the evaluation of $F(t_{n-1}+\boldsymbol{c}\,h,\boldsymbol{Y_{n}^{[k-1]}})$ in (25) and (27) (which is equivalent to the evaluation of $f(t_{n-1}+c_{i}h,Y_{n,i}^{[k-1]})$ for $i=1,\ldots,s$) also benefits from SIMD parallelization: In IRKGL16, parallelization remains transparent to the user, assuming the right-hand side of the system is defined in terms of arithmetic operations and elementary functions compatible with SIMD.jl.
 Such seamless SIMD parallelization is handled automatically with the help of SIMD.jl and an additional parametrized abstract type, VecArray{s,T}, that we have implemented to efficiently represent and handle abstract arrays of elements of type Vec{s,T} (where T is either Float64 or Float32).
-The $s$-vectors $b,c,\mu_{i},\nu_{i}$, $Y_{n}^{j}$, $F_{n}^{j}$, and $L_{n}^{j}$ considered in Subsection 2.4 are represented in IRKGL16-SIMD as vectors of type Vec{s,T}, while $Y_{n}=(Y_{n}^{1},\ldots,Y_{n}^{D})$, $F_{n}=(F_{n}^{1},\ldots,F_{n}^{D})$, and $L_{n}=(L_{n}^{1},\ldots,L_{n}^{D})$, and also $\mu=(\mu_{1},\ldots,\mu_{s})$ and $\nu=(\nu_{1},\ldots,\nu_{s})$, are represented as one-dimensional abstract arrays of type VecArray{s,T}.
-Let us consider for instance the VecArray object Yn representing $Y_{n}$:
-  • The syntax Yn[j] (or equivalently getindex(Yn,j)) can be used to get the vector of type Vec{s,T} representing $Y_{n}^{j}$.
-  • The syntax Yn[j] = V (or equivalently setindex!(Yn,j,V)) can be used to update the $j$th vector of type Vec{s,T} representing $Y_{n}^{j}$.
+The $s$-vectors $\boldsymbol{b,c,\mu_{i},\eta_{i},\nu_{i}}$, $\boldsymbol{Y_{n}^{j}}$, $\boldsymbol{F_{n}^{j}}$, and $\boldsymbol{L_{n}^{j}}$ considered in Subsection 2.5 are represented in IRKGL16 as vectors of type Vec{s,T}, while $\boldsymbol{Y_{n}=(Y_{n}^{1},\ldots,Y_{n}^{D})}$, $\boldsymbol{F_{n}=(F_{n}^{1},\ldots,F_{n}^{D})}$, and $\boldsymbol{L_{n}=(L_{n}^{1},\ldots,L_{n}^{D})}$, and also $\boldsymbol{\mu=(\mu_{1},\ldots,\mu_{s})}$, $\boldsymbol{\eta=(\eta_{1},\ldots,\eta_{s})}$, and $\boldsymbol{\nu=(\nu_{1},\ldots,\nu_{s})}$, are represented as one-dimensional abstract arrays of type VecArray{s,T}.
+Let us consider for instance the VecArray object Yn representing $\boldsymbol{Y_{n}}$:
+  • The syntax Yn[j] (or equivalently getindex(Yn,j)) can be used to get the vector of type Vec{s,T} representing $\boldsymbol{Y_{n}^{j}}$.
+  • The syntax Yn[j] = V (or equivalently setindex!(Yn,j,V)) can be used to update the $j$th vector of type Vec{s,T} representing $\boldsymbol{Y_{n}^{j}}$.
 Here, we are assuming that V is a vector of type Vec{s,T}.
@@
   • The $s$ components of Yn[j] are stored in the $j$th column of that 2-dimensional array.
-Listing LABEL:code:ex1 shows the section of the IRKGL16-SIMD code where a single vectorized fixed-point iteration (14) is implemented.
+Listing LABEL:code:ex1 shows the section of the IRKGL16 code where a single vectorized fixed-point iteration (25) is implemented.
 1f_ODE!(Fn,Yn,parms,tn+h*c)
@@
 Specifically, the function call f_ODE!(Fn, Yn, parms, tn + h*c) in Listing LABEL:code:ex1, where Fn and Yn are abstract arrays of type VecArray{s, T} (with T = Float64 or T = Float32) and tn + h*c is of type Vec{s, T}, is equivalent to the first line of (25), representing the parallel evaluation of $f(t_{n-1}+c_{i}h,Y_{n,i}^{[k-1]})$ for $i=1,\ldots,s$.
-This design enables parallel evaluations of the ODE system’s right-hand side in a way that is transparent to the user, provided that the right-hand side $f(t,y)$ of the ODE system is defined as a generic function f_ODE!(dy, y, parms, t) without branching in terms of arithmetic operations and elementary functions compatible with SIMD.jl.
-This requirement is met for all the problems considered in the numerical experiments in Section 5 below.
-However, not all ODE systems (1) support a generic implementation of f_ODE!(dy, y, parms, t) that enables seamless $s$-fold parallel evaluations of the system’s right-hand side.
-For instance, this may not be feasible if the definition of $f(t,y)$ involves unavoidable branching, which disrupts the uniformity needed for SIMD processing.
-In such cases, however, IRKGL16-SIMD can still be applied if the user provides a specialized implementation of f_ODE!(dy, y, parms, t) for instances where dy and y are abstract arrays of type VecArray{s, T} (with T = Float64 or T = Float32) and t is of type Vec{s, T}.
+This design enables parallel evaluations of the ODE system’s right-hand side in a manner that is transparent to the user, provided that the right-hand side $f(t,y)$ is implemented as a generic function f_ODE!(dy, y, parms, t) without branching and using only arithmetic operations and elementary functions compatible with SIMD.jl.
+All test problems considered in Section 5 satisfy this requirement.
+However, not all ODE systems (2) admit a generic implementation of f_ODE!(dy, y, parms, t) that supports seamless $s$-fold parallel evaluation of the system’s right-hand side.
+In particular, this may fail when $f(t,y)$ contains unavoidable branching, which breaks the uniform instruction flow required for SIMD processing.
+In such cases, the right-hand side must be evaluated sequentially for each stage vector $Y_{i}$, although the remaining vectorized computations in (24)–(25) (for first-order ODEs), (26)–(27) (for second-order ODEs) and (28) remain applicable.
+In the current implementation of IRKGL16, simultaneous SIMD evaluation of the user-supplied function $f$ is performed only when the integrator option fseq=false is explicitly selected by the user.
+All other computations are executed using SIMD vectorization whenever the state variables are of type Float32 (single precision) or Float64 (double precision).
+A fully sequential fallback implementation is used for state variables of other types (e.g., BigFloat) or whenever the simd option is explicitly disabled (simd=false) by the user.
 In some cases, such as the $N$-body example considered in Section 5, it is more convenient for the user to organize the state variables in a higher-dimensional array rather than a one-dimensional array.
@@
 Our implementations of the getindex and setindex!
-functions work as expected in this more general case, allowing users to apply IRKGL16-SIMD with a generic implementation of the function f_ODE!(dy, y, parms, t) when the state variables are stored in a higher-dimensional array.
+functions work as expected in this more general case, allowing users to apply IRKGL16 with a generic implementation of the function f_ODE!(dy, y, parms, t) when the state variables are stored in a higher-dimensional array.
 4 State-of-the-art explicit symplectic integrators
@@
 In the context of Hamiltonian systems that can be decomposed into two (or more) exactly solvable components, explicit symplectic integrators can be constructed via operator splitting techniques.
-A widely used approach is the second-order Strang splitting method [21], which yields a symmetric and symplectic scheme by composing the flows associated with the individual sub-Hamiltonians.
+A widely used approach is the second-order Strang splitting method [24], which yields a symmetric and symplectic scheme by composing the $t$-flow maps $\varphi^{A}_{t}$ and $\varphi^{B}_{t}$ associated with the individual sub-Hamiltonians.
 Specifically, the Strang splitting scheme advances the solution over a time step $h$ via the composition
-$\phi_{h}=\varphi^{A}_{h/2}\circ\varphi^{B}_{h}\circ\varphi^{A}_{h/2}.$ (21)
+$\phi_{h}=\varphi^{A}_{h/2}*\varphi^{B}_{h}*\varphi^{A}_{h/2}.$ (31)
 This method is second-order accurate, time-reversible, and symplectic.
-The resulting integrator is explicit when the Hamiltonian is written as $H=H_{A}+H_{B}$, and the exact flows $\varphi^{A}_{t}$ and $\varphi^{B}_{t}$ generated by $H_{A}$ and $H_{B}$, respectively, are available in closed form.
+The resulting integrator is explicit when the Hamiltonian is written as $H=H_{A}+H_{B}$, and the $t$-flows $\varphi^{A}_{t}$ and $\varphi^{B}_{t}$ generated by $H_{A}$ and $H_{B}$, respectively, are available in closed form.
 For instance, consider a separable Hamiltonian of the form (8), where $H_{A}(q,p)=\tfrac{1}{2}p^{T}M^{-1}p$ and $H_{B}(q,p)=U(q)$.
-In this case, the exact flows are given by $\varphi^{A}_{t}(q,p)=(q+tM^{-1}p,\,p)$ and $\varphi^{B}_{t}(q,p)=(q,\,p-t\nabla U(q))$.
+Via the linear change of variables $p=Mv$, the exact flows are given by
+$\varphi^{A}_{t}(q,v)=(q+t\,v),\quad\varphi^{B}_{t}(q,v)=(q,\,v-t\,g(q)),$ (32)
+where $g(q)=M^{-1}\nabla U(q)$.
 The method can be extended to systems where the Hamiltonian splits into more than two exactly solvable components.
 For instance, if $H=H_{A}+H_{B}+H_{C}$ then a symmetric second-order integrator can be obtained by composing the flows of each part as
-$\phi_{h}=\varphi^{A}_{h/2}\circ\varphi^{B}_{h/2}\circ\varphi^{C}_{h}\circ\varphi^{B}_{h/2}\circ\varphi^{A}_{h/2}.$ (22)
+$\phi_{h}=\varphi^{A}_{h/2}*\varphi^{B}_{h/2}*\varphi^{C}_{h}*\varphi^{B}_{h/2}*\varphi^{A}_{h/2}.$ (33)
 This generalized Strang-type splitting remains symplectic and time-symmetric, provided that each subflow $\varphi^{A}_{t}$, $\varphi^{B}_{t}$, and $\varphi^{C}_{t}$ is symplectic.
@@
 Extensions to higher-order accuracy can be obtained by symmetric compositions of $s$ basic Strang stages  [8, 13] of the form,
-$\psi_{h}=\phi_{\gamma_{s}h}\circ\phi_{\gamma_{s-1}h}\circ\dots\circ\phi_{\gamma_{2}h}\circ\phi_{\gamma_{1}h}$ (23)
+$\psi_{h}=\phi_{\gamma_{s}h}*\phi_{\gamma_{s-1}h}*\dots*\phi_{\gamma_{2}h}*\phi_{\gamma_{1}h}$ (34)
 with $\gamma_{s+1-j}=\gamma_{j}$, $1\leqslant j\leqslant s$.
-We consider the following symmetric compositions of time-symmetric second-order schemes of order $r=2,4,6,8,10$ recommended in [7]:
+We consider the following symmetric compositions of time-symmetric second-order schemes of orders $r=2,4,6,8,10$, as recommended in [8]:
   • 1)
-    Strang splitting: $r=2$th order with $s=1$ stage.
+    Strang splitting: a second-order method ($r=2$) with $s=1$ stage.
   • 2)
-    SUZ90: $r=4$th order with $s=5$ stages proposed by Suzuki(1990) [22].
+    SUZ90: a fourth-order method ($r=4$) with $s=5$ stages, proposed by Suzuki (1990) [26].
   • 3)
-    SS05: a sixth-order scheme with $s=13$ stages, a eighth-order one with $s=21$ and a tenth-order one with $s=35$ proposed by Sofroniu and Spalleta(2005) [20].
+    SS05: a sixth-order scheme with $s=13$ stages, an eighth-order one with $s=21$ stages, and a tenth-order one with $s=35$ stages proposed by Sofroniou and Spalleta (2005) [23].
 Any composition of the form (34) with (31) can be rewritten as
-$\varphi_{a_{s+1}h}^{A}\varphi_{b_{s}h}^{B}\circ\varphi_{a_{s}h}^{A}\circ\dots\circ\varphi_{b_{1}h}^{B}\circ\varphi_{a_{1}h}^{[A]}$ (24)
+$\varphi_{a_{s+1}h}^{A}\varphi_{b_{s}h}^{B}*\varphi_{a_{s}h}^{A}*\dots*\varphi_{b_{1}h}^{B}*\varphi_{a_{1}h}^{[A]}$ (35)
 with coefficients
@@
 Such more general family of methods are of particular interest in the case of Hamiltonians of the form (8), for which efficient high-order schemes have been constructed.
-They are often referred as splitting methods of Runge-Kutta-Nyström (RKN) type, as they can be rewritten as explicit symplectic RKN methods applied to the system of second order ODEs (16[18].
+They are often referred as splitting methods of Runge-Kutta-Nyström (RKN) type, as they can be rewritten as explicit symplectic RKN methods applied to the system of second order ODEs (10[21].
 We consider the following splitting integrators of RKN type of order $r=6$ and $r=8$ recommended in [8]:
@@
 5 Numerical experiments
-We compare the performance of IRKGL16-SIMD with several state-of-the-art explicit symplectic integrators for the high-precision integration of non-stiff Hamiltonian systems using constant step size.
+We compare the performance of IRKGL16 with several state-of-the-art explicit symplectic integrators for the high-precision integration of non-stiff Hamiltonian systems using constant step size.
 A range of representative examples is considered to assess both accuracy and efficiency.
-Some of the explicit symplectic integrators used in the comparison are available through the Julia language’s DifferentialEquations.jl suite [16].
-However, for consistency and fair benchmarking, we have implemented all the reference integrators described in the previous section.
-We verified that, for the problems considered, our implementations of the splitting methods are at least as efficient as—and often more so than—those available in DifferentialEquations.jl.
-The source code for the integrators, along with all Jupyter notebooks used to generate the numerical results, is publicly available at: https://github.com/mikelehu/Implicit_symplectic_can_outperform_explicit_symplectic.
-All numerical experiments were conducted on an 11th Gen Intel Core i7-11850H processor (2.5 GHz, 16 threads), a CPU equipped with 512-bit SIMD registers.
-We used the long-term support (LTS) stable release of Julia version 1.10.7 (November 26, 2024) for all tests.
-To optimize SIMD performance, Julia was launched from the terminal with bounds checking disabled using the flag --check-bounds=no.
+It is well known [13] that, for long-term integrations, it is advisable to apply some form of improved summation (e.g., Kahan’s compensated summation [15]) in (17) to prevent excessive accumulation of roundoff errors.
+The idea is to increase the accuracy of the sum by representing each component of $y_{n}$ (for $n\geqslant 0$) as the sum of two floating-point numbers while representing each component of the update $\sum_{i=1}^{n}L_{n,i}$ (expected to be of relatively smaller size) as a floating-point number.
+In IRKGL16, we employ Julia’s Base.TwicePrecision for this purpose, which provides greater accuracy (though at higher computational cost) than Kahan’s compensated summation.
+Some of the explicit symplectic integrators used in our comparison are available through the Julia DifferentialEquations.jl suite [19], but no compensated summation is applied there.
+For consistency and fair benchmarking, we implemented in Julia all the reference integrators described in the previous section using (optionally) Kahan’s compensated summation.
+We verified that, for the problems considered, our implementations of the splitting methods are at least as efficient as (and often more efficient than) the implementations in DifferentialEquations.jl when both are used without compensated summation.
+The source code for the integrators, as well as all Jupyter notebooks used to reproduce the numerical results, is publicly available in the repository associated with this article [3].
+All numerical experiments were conducted on an 11th Gen Intel Core i7-11850H processor (2.5 GHz, 16 threads) equipped with 512-bit SIMD registers.
+We used the long-term support (LTS) stable release of Julia v1.10.10 (June 27, 2025) for all tests.
 5.1 Test problems
-We have made comparisons with systems of first-order differential equations as well as with systems of second-order differential equations.
+We have made comparisons for a Hamiltonian system of first-order differential equations split into three solvable parts, and two Hamiltonian systems of second-order differential equations with Hamiltonian (8) separable in two solvable parts.
 5.1.1 System of first-order differential equations
@@
 $H_{C}(r,p_{r})$ $=-\frac{1}{r}p_{r}^{2}.$
-Although [15] provides analytical expressions for the flows $\varphi_{t}^{A}$, $\varphi_{t}^{B}$, and $\varphi_{t}^{C}$, their direct implementation is computationally inefficient.
-To enable a fair and representative comparison with our implicit symplectic integrator (IRKGL16), we have improved the evaluation of these flows, significantly enhancing the performance of the explicit method.
-The derivation and implementation details of this optimized explicit integrator are presented in Appendix A.
+Although [18] provides analytical expressions for the $t$-flow maps $\varphi_{t}^{A}$, $\varphi_{t}^{B}$, and $\varphi_{t}^{C}$, their direct implementation is computationally inefficient.
+To enable a fair and representative comparison with our implicit symplectic integrator (IRKGL16), we have improved the evaluation of these $t$-flows, significantly enhancing the performance of the explicit method.
+Unlike the analytical expressions provided in [18] for the $t$-flows $\varphi_{t}^{A}$, $\varphi_{t}^{B}$, and $\varphi_{t}^{C}$, our implementation avoids the use of trigonometric and inverse trigonometric functions entirely, resulting in improved computational efficiency.
+The derivation and implementation of the optimized flows are provided in the documentation available in the software repository [3] associated with this article.
 As in [18], we set the parameters to $E=0.995$, $L=4.6$, and $\beta=8.9\times 10^{-4}$.
@@
 The initial value of $p_{\theta}>0$ is determined by imposing the energy condition $H=-\tfrac{1}{2}$.
+The integration is performed over the interval $0,10^{5}]$.
 5.1.2 Systems of second-order differential equations
@@
 6-body outer Solar System.
-We examine a simplified model of the outer Solar System, which includes the Sun, the four outer planets (Jupiter, Saturn, Uranus, and Neptune), and Pluto, interacting via mutual Newtonian gravity.
+We examine a simplified model of the outer Solar System, which includes the Sun, the four outer planets (Jupiter, Saturn, Uranus, and Neptune), and Pluto, as point masses interacting via mutual Newtonian gravity.
 The system is described by an 18-degree-of-freedom Hamiltonian, with $q_{i},p_{i}\in\mathbb{R}^{3}$ for $i=1,\dots,6$, and Hamiltonian function
@@
 Specifically, we choose $q_{1}(0)=0$, $q_{2}(0)=0.3$, $p_{2}(0)=0.2$, and determine $p_{1}(0)>0$ such that $H=1/12$ (see [13, 14]).
-5.2 Results of Numerical Integrations
-5.2.1 Systems of First-Order Differential Equations
-Figure 1 presents work-precision diagrams for the Schwarzschild black hole problem. The left panel shows the maximum local error in the Hamiltonian,
-$\Delta H^{\text{loc}}_{\max}=\max_{n}\left(\Delta H^{\text{loc}}_{n}\right),\quad\Delta H^{\text{loc}}_{n}=\left|\frac{H(y_{n})-H(y_{n-1})}{H(y_{n-1})}\right|,$ (28)
-as a function of the total number of function evaluations.
-The right panel plots the same error against CPU time.
-In this plot, the minimal CPU time for the IRKGL16-SIMD method is highlighted with a red vertical dashed line.
-We observe that even the sequential implementation of IRKGL16 outperforms the explicit symplectic integrators for the high precision integration of this problem.
-As for the SIMD implementation, for the same CPU time budget, it achieves significantly higher accuracy even for lower precision integrations.
-Figure 1: Work-precision diagrams (maximum local error in the Hamiltonian) for the Schwarzschild black hole problem using the IRKGL16 method and symmetric composition schemes of orders $r=2,4,6,8$, and $10$.
-In the right plot, the sequential IRKGL16 implementation is denoted as IRKGL16-SEQ
-Figure 2 shows the evolution of error over time for different integrators.
-Each method is integrated using a step size chosen such that the corresponding CPU time approximately matches the red vertical dashed line in the right panel of Figure 1.
-The left panel displays the global error in the Hamiltonian,
-$\Delta H^{\text{glob}}_{\max}=\max_{n}\left(\Delta H^{\text{glob}}_{n}\right),\quad\Delta H^{\text{glob}}_{n}=\left|\frac{H(y_{n})-H(y_{0})}{H(y_{0})}\right|,$ (29)
-while the right panel shows the relative error in the radial coordinate:
-$\Delta r_{n}=\left|\frac{r_{n}-\tilde{r}_{n}}{\tilde{r}_{n}}\right|,$
-where $\tilde{r}_{n}$ denotes the reference solution obtained using IRKGL16-SIMD with a comparatively small step size ($h=16$).
+5.2 Efficiency comparison of different implementations of $s$-stage IRKGL schemes
+In Figure 1 we show work–precision diagrams for several implementations of $s$-stage IRKGL schemes applied to the Schwarzschild black hole problem.
+Figure 1: (a) Maximum local Hamiltonian error versus total function evaluations for the $s$-stage IRKGL schemes ($s=2,4,6,8$: IRKGL4, IRKGL8, IRKGL12, IRKGL16).
+(b) The same error versus CPU time for two implementations of each scheme: fully vectorized (simd=true, fseq=false) and fully sequential (simd=false, fseq=true)
+The left panel of each diagram reports the maximum local error in the Hamiltonian,
+$\Delta H^{\mathrm{loc}}_{\max}=\max_{n}\left(\Delta H^{\mathrm{loc}}_{n}\right),\qquad\Delta H^{\mathrm{loc}}_{n}=\left|\frac{H(y_{n})-H(y_{n-1})}{H(y_{n-1})}\right|,$ (39)
+as a function of the total number of function evaluations, for our implementations of the $s$-stage IRKGL schemes with $s=2,4,6,8$ (IRKGL4, IRKGL8, IRKGL12, and IRKGL16, respectively).
+This provides a measure of the intrinsic relative efficiency of the methods, independent of the computational environment, under the assumptions that (i) the user-supplied function that evaluates the right-hand side of the ODE system is executed once per stage vector, in sequential order, and (ii) the total CPU time is dominated by these function evaluations.
+According to this metric, IRKGL16 is the most efficient method in the high-accuracy regime, whereas IRKGL12 performs slightly better than IRKGL16 at moderate accuracy levels.
+The right panel plots the same error against actual CPU time for two versions of each $s$-stage IRKGL method: the fully vectorized implementation (simd=true, fseq=false) and the fully sequential one (simd=false, fseq=true).
+The relative performance of the fully sequential versions is consistent with the behavior observed in the left panels.
+In contrast, the fully vectorized implementation of IRKGL16 outperforms the lower-order schemes even at moderate accuracy.
+This improvement is due to the higher SIMD speed-up achieved by the 8-stage scheme.
+Figure 2: Panels (a)–(b) correspond to the 6-body outer Solar System, and panels (c)–(d) to the Hénon–Heiles system.
+(a, c) Maximum local Hamiltonian error versus total function evaluations for the $s$-stage IRKGL schemes ($s=2,4,6,8$).
+(b, d) The same error versus CPU time for two implementations: fully vectorized and fully sequential
+In Figure 2, we present analogous work–precision diagrams for the two second-order Hamiltonian problems.
+The results again indicate that the fully vectorized implementation of IRKGL16 is the most efficient integrator for both moderate and high accuracy levels.
+5.3 Numerical comparisons of IRKGL16 with explicit symplectic integrators
+5.3.1 Work–precision diagrams
+Figure 3: Work-precision diagrams (maximum local Hamiltonian error versus number of evaluations of $g(t,u)$) for the 6-body outer Solar System and the Hénon–Heiles problem, comparing IRKGL16 with symmetric composition schemes (orders $r=2,4,6,8,10$) and symplectic RKN methods of orders 6 and 8
+We now present work–precision diagrams for the two second-order Hamiltonian problems considered in Subsection 5.1.2.
+The computational cost of one step of any splitting method (35) applied with the flows (32) is dominated by the $s$ applications of the $t$-flow $\varphi^{B}_{t}(q,v)\mapsto(q,v+t\,g(q))$, where $g(q)=M^{-1}\nabla U(q)$.
+Thus, for splitting methods, the relative cost of a numerical integration can essentially be measured by the number of calls to the function $g(u)$.
+For implicit methods applied to Hamiltonian second-order ODEs of the form (9), the relative cost can also be quantified in terms of function evaluations of the right-hand side $g(q)$, provided that it is sufficiently expensive to compute.
+Figure 3 presents work–precision diagrams for the two second-order ODE problems, displaying the maximum local error in the Hamiltonian (39) plotted against the number of function evaluations for several splitting methods and IRKGL16.
+We consider the splitting integrators obtained by applying the composition schemes of the previous subsection with (31) as a basic integrator, and additionally the two RKN-type splitting methods of orders 6 and 8 selected in Section 4, all implemented with compensated summation.
+From the efficiency diagrams shown in Figure 3, the larger overhead inherent to implicit schemes implies that the best explicit symplectic methods will outperform sequential implementations of IRKGL16 at low and moderate accuracy levels, regardless of the programming language or other aspects of the computational environment.
+However, at higher target accuracies, i.e., when the maximum local energy error is just above roundoff level, a sequential implementation of IRKGL16 may surpass explicit methods, provided that its overhead remains sufficiently small.
+This situation frequently arises in implementations of interpreted programming languages such as MATLAB: although the user-defined evaluation of the ODE right-hand side is performed in MATLAB, the linear-algebra operations responsible for most of the overhead are executed by highly optimized underlying routines [19, 12].
+However, in other programming environments such as the Julia language, the work–precision diagrams based on function evaluations may not reflect the true computational efficiency.To quantify this, we introduce the CPU time per function evaluation ratio (CPU-FCN ratio for short), defined as the ratio of the total CPU time required for a given integration to the total number of function evaluations performed during that integration.
+For explicit splitting methods, this ratio is essentially constant: it depends neither on the step size nor on the particular scheme (i.e., the number of flows evaluated per step), since the computational cost is entirely dominated by the function evaluations themselves.
+For IRKGL16, by contrast, each step involves three additional sources of overhead beyond the function evaluations: the initial extrapolation to provide a starting guess for the fixed-point iteration, the overhead of the fixed-point iterations themselves, and the updating of state variables at the end of the step.
+Table 1 reports, for each of the second-order ODE problems, the CPU-FCN ratio of IRKGL16 normalized by that of the explicit splitting methods, for both the fully sequential implementation IRKGL16-SEQ (simd=false, fseq=true) and the fully vectorized implementation IRKGL16-SIMD (simd=true, fseq=false).
+For IRKGL16-SEQ, the normalized ratio exceeds two in both test problems, indicating that the cumulative overhead surpasses the cost of the function evaluations themselves.
+For IRKGL16-SIMD, by contrast, the normalized ratio is well below one.
+This is because the SIMD vectorization not only reduces the overhead of the implicit solver, but also enables the 8 stage function evaluations within each fixed-point iteration to be executed in parallel across SIMD registers, effectively amortizing their cost.
+As a consequence, each function evaluation in IRKGL16-SIMD is significantly cheaper in terms of wall-clock time than in the explicit methods, where function evaluations are performed sequentially.
+This implies that, relative to the work–precision diagrams presented above, the true efficiency curve of IRKGL16-SIMD is shifted to the left, making it even more competitive against the explicit symplectic integrators than those diagrams suggest.
+Table 1: CPU-FCN ratio of IRKGL16 normalized by that of the explicit splitting methods, for the two second-order ODE test problems.
+“6-body Solar System (vec)” denotes the SIMD-vectorized implementation of the flows
+IRKGL16-SEQ IRKGL16-SIMD
+Hénon–Heiles 2.5 0.6
+6-body Solar System 2.6 0.6
+6-body Solar System (vec) 3.5 0.8
+It should be noted that the CPU-FCN ratio of IRKGL16 is not constant across step sizes.
+While the overhead of the fixed-point iterations scales proportionally with the number of function evaluations per step, the overhead of the initial extrapolation and the state update does not.
+Consequently, for smaller step sizes—which require fewer fixed-point iterations per step—these fixed costs carry more relative weight, leading to a slightly larger CPU-FCN ratio.
+The values reported in Table 1 are computed at a reference step size for which the relative local energy error is approximately $10^{-14}$.
+It is worth noting that explicit splitting methods can also benefit from SIMD vectorization, by explicitly exploiting SIMD registers in the implementation of the individual flows.
+To illustrate this, we have implemented a SIMD-optimized version of the flows for the 6-body outer Solar System problem, which results in a reduced CPU-FCN ratio for the explicit methods.
+We remark that SIMD-optimizing the flows requires additional implementation effort from the user, whereas the SIMD vectorization in IRKGL16 is handled internally with a generic implementation of the user-defined function $g$.
+To assess the impact of this optimization on the relative computational cost of IRKGL16, Table 1 also reports (last row) the CPU-FCN ratio of IRKGL16 normalized by that of the SIMD-optimized explicit implementation.
+As expected, the normalized ratios increase with respect to the non-vectorized case, reflecting the fact that the explicit methods now perform each function evaluation more cheaply.
+Nevertheless, IRKGL16-SIMD retains a normalized ratio well below one, confirming that its fully vectorized implementation remains competitive even when compared against SIMD-optimized explicit integrators.
+Figure 4: Work-precision diagrams (maximum local Hamiltonian error versus CPU time) for the Schwarzschild black hole problem using symmetric composition schemes of orders $r=2,4,6,8$, and $10$, and two variants of IRKGL16: (i) fully vectorized (IRKGL16-SIMD), (ii) fully sequential (IRKGL16-SEQ)
+Figure 4 presents work–precision diagrams for the Schwarzschild black hole problem, displaying the maximum local error in the Hamiltonian (39) plotted against CPU time for explicit symplectic methods of orders $2,4,6,8,10$ and for two variants of IRKGL16: (i) the fully vectorized implementation IRKGL16-SIMD (simd=true, fseq=false), and (ii) the fully sequential implementation IRKGL16-SEQ (simd=false, fseq=true).
+In this example, we plot directly against CPU time rather than function evaluations, since the evaluation of the $t$-flows required to compute the basic map $\phi_{h}$ in (33) is considerably more expensive than the evaluation of the right-hand side of the ODE system, making the number of function evaluations a poor proxy for computational cost.
+We observe that IRKGL16 outperforms the explicit symplectic integrators across the entire accuracy range shown, with IRKGL16-SEQ already surpassing them in the high-accuracy regime.
+A key factor is that the $t$-flow evaluations in (33) are considerably more expensive than the right-hand side evaluations in IRKGL16.
+As a result, for any given CPU-time budget, IRKGL16-SIMD delivers substantially higher accuracy than the explicit methods, regardless of the accuracy level they achieve.
+5.3.2 Evolution of errors for the high-precision regime
 Figure 5: Left: global error in the Hamiltonian.
 Right: relative error in the radial coordinate $r$ for the Schwarzschild black hole problem.
-Results are shown for IRKGL16-SIMD and symmetric composition methods of orders $r=2,4,6,8,10$.
-5.2.2 Systems of Second-Order Differential Equations
-Figure 3 shows work-precision diagrams for the two second-order Hamiltonian problems.
-The top row corresponds to the 6-body outer Solar System model, while the bottom row pertains to the Hénon-Heiles system.
-The left panels display the maximum local Hamiltonian error (28) versus function evaluations, and the right panels show the same error versus CPU time.
-The red dashed line marks the optimal CPU time for the IRKGL16-SIMD method, in the sense that it is the minimal time required to achieve a local accuracy close to round-off error levels.
-Figure 3: Work-precision diagrams (maximum local Hamiltonian error) for the 6-body outer Solar System and the Hénon-Heiles problem using IRKGL16, symmetric compositions (orders $r=2,4,6,8,10$), and symplectic RKN methods of orders 6 and 8.
-Figures 46 show the evolution of errors, using simulations performed at optimal CPU time.
-The left panel in Figure 4 corresponds to the global Hamiltonian error of the 6-body outer Solar System, and the right panel to the Hénon-Heiles problem.
-Figure 4: Time evolution of the global Hamiltonian error for the 6-body outer Solar System (left) and the Hénon-Heiles problem (right), using IRKGL16, symmetric compositions, and RKN methods.
-Figure 5 depicts the position error of each planet in the outer Solar System over time, comparing IRKGL16 and SS05 ($r=10$).
-Figure 6 shows the error in both position and momentum for the Hénon-Heiles system across the different integrators.
-Figure 5: Evolution of position error for each planet in the 6-body outer Solar System using IRKGL16 and SS05 ($r=10$).
-Figure 6: Evolution of error in position (left) and momentum (right) for the Hénon-Heiles problem using IRKGL16, symmetric compositions, and RKN methods.
+Results are shown for IRKGL16-SIMD and symmetric composition methods of orders $r=2,4,6,8,10$
+We now examine the time evolution of both the global Hamiltonian error and the position error for each of the three test problems in the high-precision regime.
+For each problem, we select a reference step size $h_{\mathrm{ref}}$ for IRKGL16-SIMD such that the maximum global Hamiltonian error
+$\Delta H^{\text{glob}}_{\max}=\max_{n}\left(\Delta H^{\text{glob}}_{n}\right),\quad\Delta H^{\text{glob}}_{n}=\left|\frac{H(y_{n})-H(y_{0})}{H(y_{0})}\right|,$ (40)
+is approximately $10^{-14}$.
+Each explicit method is then run with a step size chosen so that its total CPU time matches that of IRKGL16-SIMD at $h_{\mathrm{ref}}$.
+This ensures a fair comparison on equal computational budget grounds.
+For the Schwarzschild black hole problem, the results are shown in Figure 5.
+The left panel displays the global Hamiltonian error and the right panel the relative error in the radial coordinate $r$, defined as $\Delta r_{n}=|r_{n}-\tilde{r}_{n}|/|\tilde{r}_{n}|$, where $\tilde{r}_{n}$ denotes the reference solution computed with IRKGL16-SIMD at a smaller step size ($h=h_{\mathrm{ref}}/4$).
+We compare IRKGL16-SIMD against symmetric composition methods of orders $r=2,4,6,8,10$.
+At the CPU-time-matched step sizes, the most competitive explicit methods are those of orders 2 and 4, yet their global Hamiltonian and position errors remain approximately nine orders of magnitude larger than those of IRKGL16-SIMD throughout the integration.
+Figure 6 presents the analogous comparison for the Hénon-Heiles problem, showing the global Hamiltonian error (left) and the absolute error in position $q$ (right), with the reference solution again provided by IRKGL16-SIMD at small step size.
+Here the most competitive explicit methods at equal CPU cost are the RKN-type integrators BM02 (order 6) and BCE22 (order 8), yet IRKGL16-SIMD maintains a substantially lower error in both the Hamiltonian and position throughout the integration interval.
+For the 6-body outer Solar System, the comparison is carried out against SIMD-optimized implementations of the two most competitive 8th-order explicit integrators, SS05r8 and BCE22, with step sizes matched to the CPU time of IRKGL16-SIMD using the SIMD-optimized evaluation of the acceleration function $g$.
+This represents a particularly challenging benchmark for IRKGL16, as the explicit methods benefit from problem-specific SIMD acceleration.
+Figure 7 shows the time evolution of the global Hamiltonian error: even under these conditions, IRKGL16-SIMD achieves lower Hamiltonian errors than both explicit competitors throughout the integration.
+The evolution of position errors for each planet is shown in Figure 8, comparing IRKGL16-SIMD (left) against SS05r8 (right), which exhibits lower position errors than BCE22; again, IRKGL16-SIMD yields lower position errors across all planets over the entire integration interval.
+Figure 6: Hénon-Heiles problem: time evolution of (i) the global Hamiltonian error (left), (ii) error in position (right) using IRKGL16, symmetric compositions, and RKN methods
+Figure 7: Time evolution of the global Hamiltonian error for the outer Solar System 6-body problem, computed with IRKGL16 and the 8th-order integrators SS05r8 and BCE22 using SIMD-vectorization of the flows.
+Figure 8: Evolution of the position error for each planet in the outer Solar System 6-body problem: IRKGL16 (left) and the 8th-order integrator SS05r8 using SIMD-vectorization of the flows (right).
+Across all three test problems considered—the Schwarzschild black hole, the Hénon-Heiles system, and the 6-body outer Solar System—IRKGL16-SIMD consistently achieves lower global Hamiltonian and position errors than the best-performing explicit symplectic integrators at equal CPU time in the high-precision regime.
+5.4 Discussion on alternative implementations
+The main source of overhead in implicit methods implemented with fixed-point iterations arises from the repeated updating of the state variables stored in the $D$-dimensional arrays $U_{i}$ (and $L_{i}$) at each iteration.
+At least two distinct approaches can be adopted to exploit SIMD registers in order to reduce this overhead.
+The first approach performs the updates of each state variable in the $D$-dimensional arrays $U_{i}$ independently, and vectorizes them component-wise, provided that the state components of each array $U_{i}$ are stored consecutively in memory.
+This is the more standard approach, and can in principle be handled automatically by the compiler, provided that the code is written appropriately.
+Specifically, the inner loops over $j=1,\dots,s$ required for the stage summations (16) or (20) must be unrolled, so that the outer loop over $k=1,\dots,D$ consists of straightforward arithmetic that the compiler can SIMD-vectorize.
+We have verified that, in our sequential implementation of IRKGL16, unrolling these inner loops and disabling bounds checking via the @inbounds macro causes the Julia compiler to apply component-wise vectorization automatically, substantially reducing the overhead.
+A further, marginal reduction can be achieved by applying, instead of @inbounds, the @turbo macro from the LoopVectorization.jl package.
+The second approach instead arranges, for each state variable, the corresponding components $U_{1},\dots,U_{s}$ consecutively in memory, and vectorizes their simultaneous update using SIMD $s$-vectors.
+This is the approach adopted in our implementation of IRKGL16, explained in detail in Subsection 2.5 and Section 3.
+We have verified that when the right-hand side function evaluations are performed sequentially, both vectorization approaches—the stage-wise vectorization of IRKGL16 with simd=true, fseq=false on one hand, and the component-wise vectorization via unrolled inner loops with the @turbo macro on the other—yield essentially the same reduction in overhead.
+The key advantage of stage-wise vectorization is that it enables the $s$ evaluations of the right-hand side function to be performed in parallel within SIMD registers, whenever it is implemented as a generic function compatible with the SIMD.jl package.
+This SIMD-parallel evaluation is handled internally by IRKGL16 and remains completely transparent to the user, who is only required to provide a standard generic implementation of the right-hand side function.
+In summary, we advocate for stage-wise vectorization —and have adopted it in IRKGL16 in particular—as a convenient and user-friendly implementation strategy for IRKGL schemes.
+To further assess the practical relevance of component-wise vectorization, we computed the CPU-FCN ratios of the Fortran 77 implementations of both the 12th-order IRKGL scheme and the explicit composition schemes provided in the gnicodes folder of the software repository accompanying [13], applied to the 6-body outer Solar System problem and compiled with the -O2 and -O3 optimization flags, the latter requesting vectorization whenever possible.
+We do not consider the -Ofast flag, since it enables aggressive optimization that silently eliminates compensated summation algorithms by treating their error-compensation steps as algebraically redundant, thereby compromising the long-term energy error propagation control that symplectic integrators rely on.
+For the explicit composition schemes, the CPU-FCN ratio normalized by that of our SIMD-optimized Julia implementation of the flows is approximately 1.0 under -O3 and 1.2 under -O2, suggesting that the compiler successfully applies SIMD acceleration under -O3.
+This also shows that similar efficiency can be achieved for explicit symplectic schemes in both Fortran and Julia, provided that the implementation is set up so that the compiler can exploit SIMD acceleration.
+For IRKGL12, by contrast, the normalized CPU-FCN ratio is approximately 1.4 under both -O2 and -O3, indicating that the compiler fails to apply vectorization even under -O3.
+The reason is that the number of stages $s$ is treated as a generic parameter in the gnicodes implementation, which prevents the inner summation loops from being unrolled.
+We have verified that setting $s=6$ statically, so that the number of terms in the summation is known at compile time, allows the compiler to unroll the loop and successfully apply component-wise vectorization under -O3, reducing the normalized CPU-FCN ratio to approximately 1.05 (compared to 1.4 under -O2).
+This confirms that automatic component-wise vectorization is also achievable in Fortran implementations of IRKGL schemes, provided the number of stages is fixed at compile time.
+However, even in this favorable case, the normalized CPU-FCN ratio remains above one, since no parallelism is exploited across the $s$ independent evaluations of the right-hand side function.
+By contrast, IRKGL16-SIMD achieves a normalized CPU-FCN ratio of approximately 0.8 in the 6-body Solar System problem (normalized against explicit schemes with SIMD-optimized right-hand-side function), by seamlessly and transparently exploiting stage-wise SIMD parallelism across all function evaluations.
+For ODE systems whose right-hand side is not compatible with SIMD.jl, an alternative strategy is to evaluate the $s$ stage functions in parallel using multithreading.
+In Julia, this can be conveniently implemented using the Threads.@threads macro.
+However, the additional synchronization overhead incurred by thread-level parallelism is only compensated when the right-hand side function is sufficiently expensive to evaluate [14].
+We have verified that, for the three test problems considered in this work, a straightforward application of Threads.@threads to parallelize the stage function evaluations results in increased total CPU time, indicating that the right-hand side functions in these examples are not costly enough to amortize the threading overhead.
 6 Conclusions
-IRKGL methods are symplectic and, in non-stiff regimes, can be efficiently implemented using fixed-point iterations.
-These methods combine symplecticity, time-reversibility, and high-order of accuracy, and they are broadly applicable to general canonical Hamiltonian systems [18].
-While optimized high-order explicit symplectic integrators are effective when the Hamiltonian can be decomposed into solvable components, they may be less efficient than IRKGL methods when high-precision integrations are required.
-The SIMD-based implementation of IRKGL16 significantly enhances performance, often outperforming specialized explicit symplectic methods even at moderate precision levels.
-For non-stiff problems with Hamiltonians of the specific form (17), optimized high-order explicit symplectic integrators remain highly effective due to their simplicity and computational efficiency.
-Although our sequential implementation of IRKGL16 is less efficient than the best available explicit symplectic methods for standard double-precision computations, the SIMD-enhanced version can surpass these methods in performance for high-accuracy integrations, even when constrained to double-precision arithmetic.
+High-order explicit symplectic integrators are effective when the Hamiltonian can be decomposed into solvable components, but IRKGL-based integrators offer a powerful alternative: they are symplectic, time-reversible, and high-order accurate, can be efficiently implemented using fixed-point iterations in non-stiff regimes, and are broadly applicable to general canonical Hamiltonian systems [21, 13] in first-order ODE form, regardless of whether the Hamiltonian can be split into solvable parts.
+In Section 2, we presented a general stage-wise vectorized formulation of IRKGL schemes for both first-order and second-order Hamiltonian systems.
+Beyond reducing computational overhead, this formulation leaves the algorithm naturally ready for stage-wise SIMD vectorization, enabling parallel evaluation of the right-hand side function across all stages whenever the function is compatible with SIMD.jl.
+For second-order systems, we introduced a reformulation of the corresponding Runge–Kutta–Nyström method that ensures exact symplecticity at the level of double-precision floating-point arithmetic, extending the approach of [1] originally developed for first-order systems to this setting.
+The numerical experiments of Section 5 confirm the effectiveness of the proposed approach.
+For Hamiltonian problems admitting a decomposition into a sum of exactly solvable components, IRKGL16 outperforms optimized high-order explicit symplectic integrators in the high-precision regime across all test problems considered.
+When the right-hand side function is compatible with SIMD.jl, this advantage is most pronounced; when it is not, the stage-wise vectorization of all remaining computations still greatly reduces overhead, and IRKGL16 may still surpass explicit methods at high-precision levels.
+Overall, our results show that IRKGL integrators, optimized with the proposed stage-wise SIMD vectorization technique, provide a powerful and general alternative to explicit symplectic methods, achieving high accuracy efficiently for non-stiff Hamiltonian systems.
 […]
@@
 They are also partially supported by the Department of Education of the Basque Government through the Consolidated Research Group MATHMODE (ITI456-22).
-Appendix A Efficient implementation of the three-part splitting method for the Schwarzschild black hole example
-The evaluation of the flows $\varphi_{t}^{A}$ and $\varphi_{t}^{B}$ becomes more efficient when performed in the variables $(x,y,p_{x},p_{y})$, which are related to the original coordinates by the canonical transformation
-$x$ $=r\cos\theta,\quad y=r\sin\theta,$
-$p_{x}$ $=\cos\theta\,p_{r}-\sin\theta\,\frac{p_{\theta}}{r},\quad p_{y}=\sin\theta\,p_{r}+\cos\theta\,\frac{p_{\theta}}{r}.$
-In these new variables, the Hamiltonians $H_{A}$ and $H_{B}$ become
-$\bar{H}_{A}(x,y)$ $=\frac{1}{2y^{2}}\left(L-\frac{\beta}{2}y^{2}\right)^{2}-\frac{1}{2}\left(1-\frac{2}{\sqrt{x^{2}+y^{2}}}\right)^{-1}E^{2},$
-$\bar{H}_{B}(p_{x},p_{y})$ $=\frac{1}{2}(p_{x}^{2}+p_{y}^{2}).$
-For the $H_{C}$ component, the equations of motion are simplified by introducing the variables $(\nu,\mu)$, defined through the non-canonical transformation
-$\nu=r\,p_{r},\quad\mu=\frac{p_{r}^{2}}{r}.$
-The corresponding equations are
-$\frac{d\nu}{dt}=-3\mu,\quad\frac{d\mu}{dt}=0.$
-The flows $\varphi_{t}^{A}$, $\varphi_{t}^{B}$, and $\varphi_{t}^{C}$—and hence the integrators (22) and (23)—can be efficiently evaluated in the variables $(x,y,p_{x},p_{y})$ as follows:
-  • 1)
-    Flow of $H_{A}$:
-    $\varphi_{t}^{A}:(x,y,p_{x},p_{y})\mapsto(x,y,p_{x}^{*},p_{y}^{*}),$
-    where
-    $p_{x}^{*}$ $=p_{x}-t\,\frac{\partial\bar{H}_{A}}{\partial x}(x,y),$
-    $p_{y}^{*}$ $=p_{y}-t\,\frac{\partial\bar{H}_{A}}{\partial y}(x,y).$
-  • 2)
-    Flow of $H_{B}$:
-    $\varphi_{t}^{B}:(x,y,p_{x},p_{y})\mapsto(x^{*},y^{*},p_{x},p_{y}),$
-    where
-    $x^{*}$ $=x+t\,p_{x},\quad y^{*}=y+t\,p_{y}.$
-  • 3)
-    Flow of $H_{C}$:
-    $\varphi_{t}^{C}:(x,y,p_{x},p_{y})\mapsto(x^{*},y^{*},p_{x}^{*},p_{y}^{*}),$
-    where
-    $r$ $=\sqrt{x^{2}+y^{2}},\quad c=\frac{x}{r},\quad s=\frac{y}{r},$
-    $p_{\theta}$ $=-y\,p_{x}+x\,p_{y},\quad\nu=x\,p_{x}+y\,p_{y},\quad\mu=\frac{\nu^{2}}{r^{3}},$
-    $\nu^{*}$ $=\nu-3t\,\mu,\quad p_{r}^{*}=(\mu\,\nu^{*})^{1/3},\quad r^{*}=\frac{\nu^{*}}{p_{r}^{*}},$
-    $x^{*}$ $=r^{*}\,c,\quad y^{*}=r^{*}\,s,$
-    $p_{x}^{*}$ $=c\,p_{r}^{*}-s\,\frac{p_{\theta}}{r^{*}},\quad p_{y}^{*}=s\,p_{r}^{*}+c\,\frac{p_{\theta}}{r^{*}}.$
-The partial derivatives of $\bar{H}_{A}(x,y)$ can be computed efficiently using the intermediate variables:
-$z$ $=y^{2},\quad r=\sqrt{x^{2}+z},\quad w=L-\frac{\beta}{2}z,\quad u=1-\frac{2}{r},$
-$W$ $=-\frac{w}{z},\quad U=-\frac{E^{2}}{2u^{2}},\quad R=\frac{U}{r^{3}},\quad Z=R+\frac{(W-\beta)W}{2},$
-$\frac{\partial\bar{H}_{A}}{\partial x}$ $=-2xR,\quad\frac{\partial\bar{H}_{A}}{\partial y}=-2yZ.$
-Unlike the analytical expressions provided in [15] for the flows $\varphi_{t}^{A}$, $\varphi_{t}^{B}$, and $\varphi_{t}^{C}$, our implementation avoids the use of trigonometric and inverse trigonometric functions entirely, resulting in improved computational efficiency.